### v7

# On Cone Restriction Estimates in Higher Dimensions

Abstract.
We revisit the Ou-Wang’s approach to the cone restriction problem via polynomial partitioning. By recasting their inductive scheme as a recursive algorithm and incorporating the nested polynomial Wolff axioms, we obtain improved bounds for cone restriction estimates in higher dimensions.
1. Introduction
1.1. Statement of results
In this paper, we obtain a modest improvement for the cone restriction problem in higher dimensions. The Fourier restriction problem is one of the central open questions in harmonic analysis. It concerns the possibility of restricting the Fourier transform of an $L^{p}$ function to a curved hypersurface. Stein [16] conjectured that, for a well-curved hypersurface such as the sphere, the paraboloid, or the cone considered in this paper, the Fourier transform admits a bounded restriction operator from $L^{p^{\prime}}$ to $L^{q^{\prime}}$ for an appropriate range of exponents $p^{\prime}$ and $q^{\prime}$.
We now state the cone restriction conjecture precisely.
Conjecture 1.1.
Let $n\geq 3$ and $\mathcal{C}$ be the truncated cone in $\mathbb{R}^{n}$ defined by
$\mathcal{C}:=\{(\xi_{1},\xi_{2},...,\xi_{n}):\xi_{1}^{2}+\xi_{1}^{3}+...+\xi_{n-1}^{2}=\xi_{n}^{2}\}$ (1.1)
Then for all $p,q$ satisfying
$p>\frac{2(n-1)}{n-2},\frac{n-2}{q^{\prime}}\geq\frac{n}{p}$ (1.2)
we have
$\|\hat{f}\|_{L^{q^{\prime}}(\mathcal{C};d\sigma)}\lesssim_{p,q}\|f\|_{L^{p^{\prime}}(\mathbb{R}^{n})},\forall f:\mathbb{R}^{n}\to\mathbb{C}$
where $d\sigma$ denotes the Riemannian measure on $\mathcal{C}$, and $p^{\prime},q^{\prime}$ are the Hölder conjugates of $p,q$.
Let $Ef$ be the associated Fourier extension operator on $\mathcal{C}$, i.e.,
$Ef(x):=\int_{2\bar{B}^{n-1}\setminus B^{n-1}}e^{i(x_{1}\xi_{1}+...+x_{n-1}\xi_{n-1}+x_{n}|\xi|)}f(\xi)\,d\xi\$
where $B^{n-1}\subset\mathbb{R}^{n-1}$ is the unit ball centered at the origin. By a short duality argument, we see that Conjecture 1.1 is equivalent to
Conjecture 1.2.
Let $n\geq 3$. For all $p,q$ satisfying (1.2), the estimates
$\|Ef\|_{L^{p}(\mathbb{R}^{n})}\lesssim_{p,q}\|f\|_{L^{q}(\mathbb{R}^{n-1})}$ (1.3)
holds for all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$
Conjecture 1.2 has been confirmed only in the cases $n=3,4,5$ , through the work of Barceló [1], Wolff [18], and Ou–Wang [15]. In higher dimensions, the best previously known range was obtained by Ou–Wang [15], whose results can be stated as follows.
Theorem 1.3 ([15], Theorem 2).
Let $n\geq 3$. Suppose the tuple $(p,q,k)$ is admissible in the sense that $2\leq q\leq p$, $2\leq k\leq n$ and
$\left\{\begin{array}[]{rcl}p>2\frac{n+2}{n},q^{\prime}\leq\frac{n-2}{n}p&\mathrm{if}&k=2\\ p>2\frac{n+k}{n+k-2},p\geq\frac{n}{\frac{2n-k-1}{2}-\frac{n-k+1}{q}}&\mathrm{if}&k\geq 3\end{array}\right$.
then the estimate (1.3) holds for all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$
This theorem yields the following corollary.
Corollary 1.4 ([15], Theorem 1).
Let $n\geq 3$. For all $p$ satisfying
$p>\left\{\begin{array}[]{rcl}4&\ \mathrm{if}&n=3\\ 2\frac{3n+1}{3n-3}&\mathrm{if}&n>3\quad\mathrm{odd}\\ 2\frac{3n}{3n-4}&\mathrm{if}&n>3\quad\mathrm{even}\end{array}\right$. (1.4)
the estimate
$\|Ef\|_{L^{p}(\mathbb{R}^{n})}\lesssim_{p}\|f\|_{L^{p}(\mathbb{R}^{n-1})}$ (1.5)
holds for all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$
We now state our main result. Given integers $n\geq 3$ and $k$ with $2\leq k\leq n-2$, define $p(n,k)$ and $q(n,k)$ by
$p(n,k):=2+\frac{12}{4n-5+2\left(k-1\right)\prod^{n-2}_{i=k}{\frac{2i}{2i+1}}}$
$\frac{1}{q(n,k)}=\frac{1}{p(n,k)}+\frac{1}{4n-8}\big(\frac{6}{p(n,k)}-1\big)$
For $q\geq 2$, we define $p_{n,k}(q)$ by
[…]
where $\alpha_{n,k}(q)$ is defined by
$\frac{1}{q}=\frac{\alpha_{n,k}(q)}{2}+\frac{1-\alpha_{n,k}(q)}{q(n,k)}$
Theorem 1.5.
Let $n\geq 3$. Suppose the tuple $(p,q,k)$ is admissible in the sense that $2\leq q\leq p$, $2\leq k\leq n-2$ and
[…]
then the estimate (1.3) holds for all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$
Arguing as in [12], we obtain
Corollary 1.6.
For $n\geq 3$ and all $f:\mathbb{R}^{n-1}\mapsto\mathbb{C}$, the estimate (1.5) holds whenever
$p>2+\lambda n^{-1}+O(n^{-2}),\quad\lambda=2.596$… (1.6)
Observe that (1.4) implies that the estimate (1.5) holds for all $f:\mathbb{R}^{n-1}\to\mathbb{C}$ whenever
$p>2+\frac{8}{3n}+O(n^{-2}).$
Consequently, Corollary 1.6 yields an improved $L^{p}\to L^{p}$ bound for the cone restriction conjecture in higher dimensions.
Let $B_{R}\subset\mathbb{R}^{n}$ denote a ball of radius $R$. By a real interpolation of restricted-type, together with a standard $\epsilon$-removal argument (see [17, 4, 21]) and a bilinear interpolation (see [15]), Theorem 1.3 reduces to
Proposition 1.7.
Let $n\geq 3$. Suppose the tuple $(p,q,k)$ satisfies $2\leq q\leq p$, $2\leq k\leq n-2$, and
$\begin{cases}p\geq p_{n,2}(q),q^{\prime}\leq\frac{n-2}{n}p&\text{if $k=2$},\\[4.0pt] p\geq p_{n,k}(q),p\geq\frac{n}{\frac{2n-k-1}{2}-\frac{n-k+1}{q}}&\text{if $k\geq 3$}.\\[4.0pt] \end{cases},$
then the estimate
$\|Ef\|_{L^{p}(B_{R})}\lesssim_{p,q,\varepsilon}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{q}}\|f\|_{L^{\infty}(\mathbb{R}^{n-1})}^{1-\frac{2}{q}}$ (1.7)
holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$.
The Fourier restriction problem has been the focus of extensive study for decades. For hypersurfaces with non-vanishing Gaussian curvature, such as the sphere and paraboloid, restriction estimates are closely tied to several central problems in harmonic analysis, discrete geometry, and PDE, including the Bochner–Riesz conjecture, the Kakeya conjecture, and the Schrödinger maximal estimate. Furthermore, Fourier restriction has applications beyond harmonic analysis, impacting fields such as geometric measure theory (e.g., Falconer’s distance set conjecture) and number theory (e.g., counting solutions to Diophantine equations). However, sharp restriction estimates are known only for very few surfaces and dimensions. For instance, the restriction conjecture for the paraboloid and sphere remains unsolved for $n\geq 3$. It is known that the restriction conjecture for paraboloids or spheres in $\mathbb{R}^{n}$ implies the corresponding conjecture for cones in $\mathbb{R}^{n+1}$ (see [14]), which highlights the potential applications of our results.
Guth [5, 6] introduced polynomial partitioning to the study of Fourier restriction and established an improved results for paraboloids. Later, his results were subsequently refined by Wang[19], Wang-Wu[20], Hickman–Rogers [11], Hickman–Zahl [12], and Guo-Wang-Zhang [7]. In [11, 12] the authors reformulated Guth’s induction-on-scale argument into a recursive algorithm, utilizing various Kakeya-type geometric results from [13, 25], including the nested polynomial Wolff axiom, to bound the number of directions of the tubes those stay in varieties. Beyond paraboloid restriction, polynomial partitioning has also been applied to many related problems in harmonic analysis, such as Bochner-Riesz conjecture (see [22, 8]), the Schrödinger maximal estimate (see [3]), and cone restriction (see [15]).
In this article, we refine the results presented in [15] to establish an improved $k$-broad estimate for the cone by adapting the methods used in [11, 12]. Specifically, we will employ iterated polynomial partitioning and the nested polynomial Wolff axiom in Section 3, where we reformulate the core of the inductive proofs from [15] as a recursive process. Compared to the paraboloid case addressed in [11, 12], the main difference in our approach is as follows: in the paraboloid case, a leaf backtracks to the root. However, if we were to apply this strategy to the cone, the number of tube directions would not be well-bounded, as the number of sectors on a cone is significantly fewer than the number of caps on a paraboloid. Instead, we make a leaf backtrack to its (n1)(n-1)- dimensional ancestor.
1.2. Structure of the article
This article is not self-contained and heavily references the work of Guth [6], Ou-Wang [15], Hickman-Rogers [11], and Hickman-Zahl [12]. In Section 2, we review some of the tools employed in this work, including the $k$-broad norm, wave packet decomposition, polynomial partitioning, transverse equidistribution estimates, and the nested polynomial Wolff axiom. In Section 3, we present a modified version of the polynomial structural decomposition described in [11, 12] by applying iterated polynomial partitioning and the nested polynomial Wolff axiom. Finally, Section 4 discusses the numerology of Theorem 1.7.
1.3. Notational Conventions
In this paper, we work with Schwartz functions $f$ and $g$ supported in $2\overline{B}^{\,n-1}\setminus B^{n-1}$. We write $x\in\mathbb{R}^{n}$ as $x=(x^{\prime},x_{n})$. The parameters $n$, $k$, and $\epsilon$ are fixed, and all implicit constants are allowed to depend on them.
We inductively define $\delta$ and $\delta_{j}$ for $0\leq j\leq n$ by
$\delta_{0}:=\epsilon^{10},\qquad\delta_{j}:=\delta_{j-1}^{10}\quad(1\leq j\leq n),\qquad\delta:=\delta_{n}^{10}.$
We write $A\lesssim_{L}B$, $B\gtrsim_{L}A$, or $A=O_{L}(B)$ if $A\leq C_{L}B$ for some constant $C_{L}$ depending only on $L$. We write $A\sim_{L}B$ if both $A\lesssim_{L}B$ and $A\gtrsim_{L}B$ hold. We write $A\ll B$ or $B\gg A$ if $CA\leq B$ for some admissible constant $C\geq 1$ that may be taken arbitrarily large. We write $A\lessapprox B$ if $A\lesssim R^{O(\delta_{0})}B$.
Let #A\#A denote the cardinality of a finite set $A$. Let $N_{r}E$ denote the $r$-neighbourhood of a set $E$. Let $B_{r}\subset\mathbb{R}^{n}$ denote an arbitrary ball of radius $r$. Let $T_{x}Z$ denote the tangent space of $Z$ at $x$.
We say that $A^{\prime}\subseteq A$ is a refinement of $A$ if $A\lessapprox\#A^{\prime}$. We write $A=\mathrm{RapDec}(r)$ if $A\lesssim_{N}r^{-N}$. We say that a function $F$ is essentially supported in $\Omega$ if
$F(x)\lesssim\mathrm{RapDec}(R)\,\|f\|_{2}\quad\text{for all }x\notin\Omega.$
1.4. Acknowledgements
The author would like to thank Xiaochun Li for suggesting the problem and for numerous insightful discussions. The author is also grateful to Larry Guth and Shunhua Shizuo for their valuable guidance in learning decoupling theory.
1.5. Funding
This research received no external funding.
2. Preliminaries
2.1. Wave Packet Decomposition
Wave packet decomposition is a central tool in our argument. We briefly recall its definition and basic properties from Sections 2.2 and 5.2.1 of [15].
Definition. Fix $R\gg 1$ and let $R^{\delta_{0}}\leq r\leq R$. Cover $2\overline{B}^{\,n-1}\setminus B^{n-1}$ by finitely overlapping $r^{-1/2}$-sectors $\theta\in\Theta[r]$. A sector $\theta\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ is called an $r^{-1/2}$-sector if it has unit length in the radial direction and angular width $r^{-1/2}$. We denote by $\Theta[r]$ the collection of all such finitely overlapping $r^{-1/2}$-sectors. Given a Schwartz function $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$, we write
$f=\sum_{\theta\in\Theta[r]}f_{\theta},\qquad f_{\theta}:=f\,\psi_{\theta},$
where $\{\psi_{\theta}\}$ is a smooth partition of unity subordinate to the cover $\{\theta\}$. For each $\theta$, we further decompose $f_{\theta}$ as follows. First, decompose $\mathbb{R}^{n-1}$ into finitely overlapping cubes of side length $r^{1/2+\delta}$ centered at $v\in r^{1/2+\delta}\mathbb{Z}^{n-1}$. Then
$\widehat{f_{\theta}}=\sum_{v}\widehat{f_{\theta}}\,\eta_{v},$
where $\{\eta_{v}\}$ is a smooth partition of unity subordinate to this cover.
Next, for each $v$, we decompose $\widehat{f_{\theta}}\eta_{v}$ into finer pieces adapted to $\theta$. Cover the cube centered at $v$ by parallel thin plates $P_{\theta,v}^{l}$ of radius $r^{1/2+\delta}$, thickness $r^{\delta/2}$, and normal direction $\xi_{\theta}$, where $\xi_{\theta}$ denotes the unit vector pointing in the central direction of the sector $\theta$. Hence,
$f=\sum_{(\theta,v,l)\in\mathbb{T}[r]}\big(\widehat{f_{\theta}}\,\eta_{v}\,\eta_{\theta,v}^{l}\big)^{\vee},$
where $\{\eta_{\theta,v}^{l}\}$ is a smooth partition of unity subordinate to this cover, and $\mathbb{T}[r]$ denotes the collection of wave packets at scale $r$.
Since $\big(\widehat{f_{\theta}}\,\eta_{v}\,\eta_{\theta,v}^{l}\big)^{\vee}$ is essentially supported in $\theta$ (more precisely in $C\theta$ for some constant $C$), we choose bump functions $\psi_{\theta}^{\prime}$ such that $\psi_{\theta}^{\prime}=1$ on an $r^{-1/2}$–neighbourhood of $\operatorname{supp}\psi_{\theta}$ and decays rapidly outside this region. We then define
$f_{\theta,v}^{l}:=\psi_{\theta}^{\prime}\big(\widehat{f_{\theta}}\,\eta_{v}\,\eta_{\theta,v}^{l}\big)^{\vee}.$
Consequently,
$f=\sum_{(\theta,v,l)\in\mathbb{T}[r]}f_{\theta,v}^{l}\;+\;\mathrm{RapDec}(R)\,\|f\|_{2}.$ (2.1)
For any $\mathcal{T}\subseteq\mathbb{T}[r]$, define
$f_{\mathcal{T}}:=\sum_{(\theta,v,l)\in\mathcal{T}}f_{\theta,v}^{l}.$
We say that $f$ is concentrated on the wave packets from $\mathcal{T}\subseteq\mathbb{T}[r]$ if
$\|f_{\mathbb{T}[r]\setminus\mathcal{T}}\|_{\infty}\lesssim\mathrm{RapDec}(R)\,\|f\|_{2}.$
$L^{2}$-Orthogonality. By Plancherel theorem, the functions $f_{\theta,v}^{l}$ are almost orthogonal in $L^{2}$.
Lemma 2.1 ([15], Section 2.2).
Suppose $r\geq 1$. For any $\mathcal{T}\subseteq\mathbb{T}[r]$, there holds
$\|f_{\mathcal{T}}\|_{2}^{2}\sim\sum_{(\theta,v,l)\in\mathcal{T}}\|f_{\theta,v}^{l}\|_{2}^{2}$
As in the paraboloid case [11], there are corresponding local versions of the $L^{2}$-orthogonality for the cone, which follow directly from Lemma 2.1:
Corollary 2.2.
Let $1\leq\rho\leq r$. For any set $\mathcal{T}\subseteq\mathbb{T}[r]$ and $\rho^{-1/2}$-sector $\theta^{*}$, we have
$\|\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{f^{l}_{\theta,v}}\|^{2}_{L^{2}\left({\theta}^{*}\right)}\lesssim\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{{\left\|f^{l}_{\theta,v}\right\|}^{2}_{L^{2}\left(C{\theta}^{*}\right)}}$
$\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{{\left\|f^{l}_{\theta,v}\right\|}^{2}_{L^{2}\left({\theta}^{*}\right)}}\lesssim{\left\|\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{f^{l}_{\theta,v}}\right\|}^{2}_{L^{2}\left(C{\theta}^{*}\right)}$
for some constant $C$.
Corollary 2.3.
Let $1\leq\rho\leq r$. For any set $\mathcal{T}\subseteq\mathbb{T}[r]$, we have
${\mathop{\mathrm{max}}_{{\theta}^{*}\mathrm{:}{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{f^{l}_{\theta,v}}\right\|}^{2}_{L^{2}\left({\theta}^{*}\right)}\ }\lesssim{\mathop{\mathrm{max}}_{{\theta}^{*}\mathrm{:}{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{{\left\|f^{l}_{\theta,v}\right\|}^{2}_{L^{2}\left(C{\theta}^{*}\right)}}\ }$
${\mathop{\mathrm{max}}_{{\theta}^{*}\mathrm{:}{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{{\left\|f^{l}_{\theta,v}\right\|}^{2}_{L^{2}\left({\theta}^{*}\right)}}\ }\lesssim{\mathop{\mathrm{max}}_{{\theta}^{*}\mathrm{:}{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|\sum_{\left(\theta,v,l\right)\in\mathcal{T}}{f^{l}_{\theta,v}}\right\|}^{2}_{L^{2}\left(C{\theta}^{*}\right)}\ }$
for some constant $C$.
Spatial Concentration. For a wave packet $(\theta,v,l)$, the restriction of $Ef_{\theta,v}^{l}$ to the ball $B(0,r)$ is essentially supported in a tube. More precisely, define
$T_{\theta,v}^{l}:=N_{r^{\delta}}\!\left(P_{\theta,v}^{l}+rL(\theta)\right),\qquad rL(\theta):=\{\,t\,L(\theta):-r\leq t\leq r\,\},$
which is an $r$-tube with long direction $L(\theta)$ and short direction $M(\theta)$, where
$L(\theta):=(-\xi_{\theta},\,1),\qquad M(\theta):=(\xi_{\theta},\,1).$
It is well known that $Ef_{\theta,v}^{l}$ is essentially supported in $T_{\theta,v}^{l}$.
2.2. A Standard $L^{2}$ Estimate.
Moreover, there is a standard $L^{2}$ estimate for EfEf; see [6, 15] for details.
Lemma 2.4.
Suppose $f$ is a Schwartz function and supported in $2{\overline{B}}^{n-1}\backslash B^{n-1}\to\mathbb{C}$, then
${\left\|Ef\right\|}_{L^{2}\left(B_{R}\right)}\lesssim R^{1/2}{\left\|f\right\|}_{2}$
Also, if $R^{1/2+\delta}\leq r\leq R$ and $f$ is concentrated on $R$-scale wave packets $\left(\theta,v,l\right)$ with $T^{l}_{\theta,v}\cap B_{r}\neq\emptyset$, then
${\left\|Ef\right\|}_{L^{2}\left(B_{10r}\right)}\sim r^{1/2}{\left\|f\right\|}_{2}$
2.3. Comparing Wave Packet Decompositions at Different Scales.
Let $r^{1/2+2\delta}\leq\rho\leq r$ and let $y\in B(0,r)$. Suppose $g$ is concentrated on an $r$-scale wave packet $(\theta,v,l)$ with $T_{\theta,v,l}\cap B(y,\rho)\neq\emptyset$. To perform induction on scales, we decompose $g$ into wave packets at the smaller scale $\rho$ inside the ball $B(y,\rho)$. We begin by recentering $B(y,\rho)$ at the origin via the modulation
$\tilde{g}(\xi):=e^{i\psi_{y}(\xi)}\,g(\xi),$
where
${\psi}_{y}\left(\xi\right):=y_{1}{\xi}_{1}+\dots+y_{n-1}{\xi}_{n-1}+y_{n}|\xi|$
[…]
$\tilde{g}=\sum_{\left(\zeta,w,L\right)\in\mathbb{T}\left[\rho\right]}{{\tilde{g}}^{L}_{\zeta,w}}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}_{2}$
Let $T^{L}_{\zeta,w}\left(y\right)\$denote the $\rho$-tube corresponding to $\left(\zeta,w,L\right)$ in the $x$ coordinate (contained in $B\left(y,\rho\right)$). Define ${\widetilde{\mathbb{T}}}_{\theta,v,l}$ as
${\widetilde{\mathbb{T}}}_{\theta,v,l}:=\left\{\left(\zeta,w,L\right):\ \mathrm{Dist}\left(\theta,\zeta\right)\lesssim{\rho}^{-1/2},\mathrm{Dist}\left(P^{L}_{\zeta,w},P^{l}_{\theta,v}+P^{0}_{\theta}-{\partial}_{\xi}{\psi}_{y}\left({\xi}_{\theta}\right)\right)\lesssim r^{\delta}\right\}$
where
$\mathrm{D}\mathrm{ist}\left(A,B\right):={\mathop{\mathrm{max}}_{a\in A}{\mathop{\mathrm{min}}_{b\in B}\mathrm{dist}\left(a,b\right)\ }\ }\ \mathrm{for}\ A,B\subseteq{\mathbb{R}}^{n}$
and $P^{0}_{\theta}$ denotes the thin plate centered at the origin, with radius $\sim r^{1/2+\delta}$, thickness $\sim r^{\delta/2}$, and normal direction $\xi_{\theta}$.
Lemma 2.5 ([15], Lemma 5.4).
For any $\left(\zeta,w,L\right)\in{\widetilde{\mathbb{T}}}_{\theta,v,l}$, we have
$\angle\left(L\left(\theta\right),L\left(\zeta\right)\right)\lesssim{\rho}^{-1/2},\angle\left(M\left(\theta\right),M\left(\zeta\right)\right)\lesssim{\rho}^{-1/2}$
$\mathrm{D}\mathrm{ist}\left(T^{L}_{\zeta,w}\left(y\right),\ \left[T^{l}_{\theta,v}\cap B\left(y,\rho\right)\right]+P^{0}_{\theta}\right)\lesssim r^{\delta}$
where $\angle\left(v_{1},v_{2}\right)$denotes the angle between two non-zero vectors $v_{1},v_{2}\in{\mathbb{R}}^{n}$.
The relationship between these two wave packet decompositions was studied in detail in [15]. In particular, if $g$ is concentrated on wave packets from $\mathcal{T}\subseteq\mathbb{T}[r]$, one may ask which wave packets at scale $\rho$ the function $\tilde{g}$ is concentrated on. We recall the relevant results from [15] below.
Lemma 2.6 ([15], Lemma 5.3).
If $g$ is concentrated on wave packet from $\mathcal{T}\subseteq\mathbb{T}\left[r\right]$, then $\tilde{g}$ is concentrated on wave packets in
$\widetilde{\mathcal{T}}=\bigcup_{\left(\theta,v,l\right)\in\mathcal{T}}{{\widetilde{\mathbb{T}}}_{\theta,v,l}}$
Roughly speaking, Lemma 2.6 asserts that $(g_{\theta,v}^{l})^{\widetilde{}}$ is concentrated on scale $\rho$ wave packets that are parallel to $T_{\theta,v}^{l}$ and contained in $c\,T_{\theta,v}^{l}$ for some constant $c$.
2.4. Medium Tubes.
Next, we group $\mathbb{T}[r]$ and $\mathbb{T}[\rho]$ into subcollections. For a $\rho^{-1/2}$-sector $\zeta_{0}$ and $v_{0}\in r^{1/2+2\delta}\mathbb{Z}^{n-1}\cap B(0,\rho)$, let $\mathbb{T}_{\zeta_{0},v_{0}}(y)$ denote the collection of $r$-tubes $T_{\theta,v}^{l}$ such that
${\mathrm{D}\mathrm{ist}}_{H}\left({\zeta}_{0},\theta\right)\lesssim{\rho}^{-1/2},\emptyset\neq T^{l}_{\theta,v}\cap B\left(y,\rho\right)\subseteq B\left(v_{0},r^{1/2+2\delta}\right)+\rho L\left({\zeta}_{0}\right)+\left\{y\right\}$
here $\mathrm{Dist}_{H}$ denotes the Hausdorff distance. If an $r$-tube $T_{\theta,v}^{l}$ satisfies the above conditions for multiple pairs $(\zeta_{0},v_{0})$, we assign $T_{\theta,v}^{l}$ to only one $\mathbb{T}_{\zeta_{0},v_{0}}(y)$. We also define
$g_{{\zeta}_{0},v_{0}}:=\sum_{\left(\theta,v,l\right)\in{\mathbb{T}}_{{\zeta}_{0},v_{0}}\left(y\right)}{g^{l}_{\theta,v}}$
${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}:=\bigcup_{\left(\theta,v,l\right)\in{\mathbb{T}}_{{\zeta}_{0},v_{0}}\left(y\right)}{{\widetilde{\mathbb{T}}}_{\theta,v,l}},\ \ {\tilde{g}}_{{\zeta}_{0},v_{0}}:=\sum_{\left(\zeta,w,L\right)\in{\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}}{{\tilde{g}}^{L}_{\zeta,w}}$
We immediately conclude that for all $(\zeta,w,L)\in{\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$
${\mathrm{Dist}}_{H}({\zeta}_{0},\zeta)\lesssim{\rho}^{-\frac{1}{2}},P^{L}_{\zeta,w}\subseteq B(v_{0},r^{\frac{1}{2}+2\delta})$
and
$g=\sum_{{\zeta}_{0},v_{0}}{g_{{\zeta}_{0},v_{0}}}+\mbox{RapDec}\left(r\right){\left\|g\right\|}_{2}$
Since each $\left(\zeta,w,L\right)\$belongs to $\sim 1$ ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$, Lemma 2.1 implies
${\left\|g\right\|}^{2}_{2}\sim\sum_{{\zeta}_{0},v_{0}}{{\left\|g_{{\zeta}_{0},v_{0}}\right\|}^{2}_{2}},\ \ {\left\|\tilde{g}\right\|}^{2}_{2}\sim\sum_{{\zeta}_{0},v_{0}}{{\left\|{\tilde{g}}_{{\zeta}_{0},v_{0}}\right\|}^{2}_{2}}$ (2.2)
It follows from Lemmas 2.5 and 2.6 that
Corollary 2.7 ([15], Lemma 5.6).
If $g$ is concentrated on wave packets in ${\mathbb{T}}_{{\zeta}_{0},v_{0}}\left(y\right)$, then $\tilde{g}$ is concentrated on wave packets in ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$. On the other hand, if $\tilde{g}$ is concentrated on wave packets in ${\widetilde{\mathbb{T}}}_{{\zeta}_{0},v_{0}}$, then inside $B\left(y,\rho\right)$$g$ is concentrated on wave packets in
$\bigcup_{\begin{array}[]{c}{\mathrm{D}\mathrm{ist}}_{H}\left({\zeta}_{0},\zeta\right)\lesssim{\rho}^{-1/2}\\ {\mathrm{D}\mathrm{ist}}_{H}\left(v_{0},v\right)\lesssim r^{1/2+2\delta}\end{array}}{{\mathbb{T}}_{\zeta,v}\left(y\right)}$
2.5. Tangential and Transverse Wave Packets.
We recall several definitions and properties of tangential and transverse wave packets from [6, 15].
Definition 2.8 (Transverse complete intersection, [6], Section 5.1).
Suppose $1\leq m\leq n$. Let $Z\left(P_{1},\dots,P_{n-m}\right)$ denote the set of common zeros of the polynomials $P_{1},\dots,P_{n-m}$ on ${\mathbb{R}}^{n}$. The variety $Z\left(P_{1},\dots,P_{n-m}\right)$ is said to be a transverse complete intersection if
$\nabla P_{1}\left(x\right)\wedge\dots\wedge\nabla P_{n-m}\left(x\right)\neq 0,\ \ \forall x\in Z\left(P_{1},\dots,P_{n-m}\right)$
The degree of the transverse complete intersection is defined as $\mathop{\mathrm{max}}_{1\leq j\leq n-m}{\mathrm{deg}P_{j}\ }\$}.
From the above definition, if $\mathrm{Z}:=Z(P_{1},\dots,P_{n-m})$ is a transverse complete intersection, then $\mathrm{Z}$ is a smooth manifold of dimension $m$. In particular, $\mathrm{Z}$ admits a tangent space $T_{x}\mathrm{Z}$ at every point $x\in\mathrm{Z}$.
Definition 2.9 (Tangential and transverse tubes, [6], Section 5.2).
Let $\bm{\mathrm{Z}}$ be a transverse complete intersection of dimension $m$. For a fixed ball $B\left(y,r\right)$, a $r$-tube $T^{l}_{\theta,v}$ with $T^{l}_{\theta,v}\cap B\left(y,r\right)\cap\bm{\mathrm{Z}}\neq\emptyset$ is said to be tangent to $\mathrm{Z}$ in $B\left(y,r\right)$ if $T^{l}_{\theta,v}\cap B\left(y,2r\right)\subseteq N_{r^{1/2+{\delta}_{m}}}\left(\bm{\mathrm{Z}}\right)$ and
$\angle\left(L\left(\theta\right),T_{x}\bm{\mathrm{Z}}\right)\lesssim r^{-1/2+{\delta}_{m}},\ \ \forall x\in\mathrm{Z}\cap B\left(y,r\right)\cap N_{r^{1/2+{\delta}_{m}}}\left(T^{l}_{\theta,v})\right)$
Otherwise, we say that $T^{l}_{\theta,v}$ is transverse to $\bm{\mathrm{Z}}$ in $B\left(y,r\right)$. We use $\mathbb{T}\left[\bm{\mathrm{Z}},B\left(y,r\right)\right]$ to denote the set of $r$-tubes tangent to $\bm{\mathrm{Z}}$ in $B\left(y,r\right)$.
In Section 3, we require the following result, which can be viewed as a continuum analog of Bézout’s theorem, to control the transverse intersections between a tube and an algebraic variety.
Lemma 2.10 ([6], Lemma 5.7).
Let $T$ be a cylinder of radius $r$ with central line $l$ and suppose that $Z=Z\left(P_{1},\dots,P_{n-m}\right)\subseteq{\mathbb{R}}^{n}$ is a transverse complete intersection with degree at most $D$. For any $\alpha>0$, define
$Z_{>\alpha}:=\left\{x\in Z:\angle\left(l,T_{x}\bm{\mathrm{Z}}\right)>\alpha\right\}$
Then $Z_{>\alpha}\cap T$ is contained in a union of $\lesssim D^{n}$ balls of radius $\lesssim r{\alpha}^{-1}$.
2.6. Reduction to Broad Estimates
In [6], Guth decomposed $\|Ef\|_{L^{p}(B_{R})}$ into a broad part and a narrow part. The narrow part is essentially supported in a lower-dimensional subspace of $\mathbb{R}^{n}$. Using decoupling [2] together with induction on scales, he reduced the linear restriction estimates to estimating the broad part. Ou–Wang [15] followed the same strategy.
Fix a large constant $K$. Decompose $2\overline{B}^{\,n-1}\setminus B^{n-1}$ in frequency space into $K^{-1}$-sectors $\tau$. Given a Schwartz function $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$, write
$f=\sum_{\tau}f_{\tau},\qquad f_{\tau}:=f\,\psi_{\tau},$
where $\{\psi_{\tau}\}_{\tau}$ is a smooth partition of unity subordinate to the cover $\{\tau\}$. Cover $B_{R}$ by finitely overlapping $B_{K^{2}}$. Fix a parameter $1<A\lesssim K^{\epsilon}$. For each $B_{K^{2}}$, define $\mu_{Ef}(B_{K^{2}})$ by
$\mu_{Ef}(B_{K^{2}}):={\mathop{\mathrm{min}}_{V_{1},\dots,\ V_{A}\ \left(k-1\right)\mathrm{-}\mathrm{subspace\ of}\ {\mathbb{R}}^{n}}\left(\mathop{\mathrm{max}}_{\tau:\angle\left(L\left(\tau\right),V_{a}\right)\gtrsim K^{-2},\forall 1\leq a\leq A}\int_{B_{K^{2}}}{{\left|Ef_{\tau}\right|}^{p}}\right)\ }$
where
$\angle\left(L\left(\tau\right),V_{a}\right):={\mathop{\mathrm{min}}_{v\in L\left(\tau\right)}\angle\left(v,V_{a}\right)\ }$
For $U\subseteq{\mathbb{R}}^{n}$ the $k$-broad norm over $U$ is defined by
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(U\right)}:={\left(\sum_{B_{K^{2}}}{\frac{\left|B_{K^{2}}\cap U\right|}{\left|B_{K^{2}}\right|}}{\mu}_{Ef}\left(B_{K^{2}}\right)\right)}^{\frac{1}{p}}$
Here we recall some basic properties about broad norm.
Lemma 2.11 (Triangle inequality, [15], Lemma 3.1).
Suppose that $1\leq p<\infty$, $f=g+h$, and $A=A_{1}+A_{2}$, where $A,A_{1},A_{2}$ are non-negative integers. Then
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(U\right)}\lesssim{\left\|Eg\right\|}_{BL^{p}_{k,A_{1}}\left(U\right)}+{\left\|Eh\right\|}_{BL^{p}_{k,A_{2}}\left(U\right)}$
Lemma 2.12 (Hölder’s inequality, [15], Lemma 3.2).
Suppose $1\leq p,p_{1},p_{2}<\infty$, and $0\leq{\alpha}_{1},{\alpha}_{2}\leq 1$ obey ${\alpha}_{1}+{\alpha}_{2}=1$ and
$\frac{1}{p}=\frac{{\alpha}_{1}}{p_{1}}+\frac{{\alpha}_{2}}{p_{2}}$
Suppose that and $A=A_{1}+A_{2}$, then
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(U\right)}\leq{\left\|Ef\right\|}^{{\alpha}_{1}}_{BL^{p_{1}}_{k,A_{1}}\left(U\right)}{\left\|Ef\right\|}^{{\alpha}_{2}}_{BL^{p_{2}}_{k,A_{2}}\left(U\right)}$
Lemma 2.13 (Standard $L^{2}$-estimate, [15], Section 5).
For $f\in L^{2}\left(2{\overline{B}}^{n-1}\backslash B^{n-1}\right)$ and non-negative integers $A$, there holds
${\left\|Ef\right\|}^{2}_{BL^{2}_{k,A}\left(B_{r}\right)}\lesssim r{\left\|f\right\|}^{2}_{2}$
Lemma 2.14 (Vanishing property, [15], Lemma 4.5).
Suppose $Ef$ is tangent to a transverse complete intersection $\bm{Z}$ in $B_{r}$. If $\mathrm{deg}\bm{Z}\lesssim 1\$} and $dim\bm{Z}\$} $\leq k-1$, then
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{r}\right)}\lesssim\mathrm{RapDec}\left(r\right){\left\|f\right\|}_{2}$
Broad estimates may imply linear estimates. More precisely,
Theorem 2.15 ([15], Section 6.2).
Let $n\geq 3$ and $2\leq q\leq p$. Suppose for some integer $A>0$ the estimate
$\|Ef\|_{BL_{k,A}^{p}(B_{R})}\lesssim_{p,q,\varepsilon}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{q}}\|f\|_{L^{\infty}(\mathbb{R}^{n-1})}^{1-\frac{2}{q}}$ (2.3)
holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$. If there exists an integer $k$ with $2\leq k\leq n$ and
$\left\{\begin{array}[]{rcl}q^{\prime}\leq\frac{n-2}{n}p&\mathrm{if}&k=2\\ p\geq\frac{n}{\frac{2n-k-1}{2}-\frac{n-k+1}{q}}&\mathrm{if}&k\geq 3\end{array}\right$.
then the estimate (1.7) holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$.
Hence, together with Lemma 2.12, Proposition 1.7 reduces to
Proposition 2.16.
Suppose $n\geq 3$ and $2\leq k\leq n-2$. For $p=p(n,k)$, $q=q(n,k)$ and some integer $A\lesssim_{\varepsilon,n}1$, the estimate (2.3) holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$.
2.7. Transverse Equidistribution Estimates
Using a version of the Heisenberg uncertainty principle, Guth [6] established transverse equidistribution estimates for the paraboloid. However, in the cone setting the tubes are much thinner, and $Eg$ may be concentrated on an extremely thin layer of $N_{R^{1/2+\delta_{m}}}(\mathbf{Z})$, so this property may fail to hold.
To overcome this geometric obstruction, [15] decomposes $g$ into an essential part $g_{\mathrm{ess}}$ and a tail part $g_{\mathrm{tail}}$. Roughly speaking, $g_{\mathrm{ess}}$ is concentrated on wave packets $(\theta,v,l)$ for which $M(\theta)$ is not perpendicular to the tangent space of $\mathbf{Z}$ at some point $z\in\mathbf{Z}\cap N_{R^{1/2+\delta_{m}}}(T_{\theta,v}^{l})$; consequently, transverse equidistribution estimates hold for $g_{\mathrm{ess}}$. The tail term $g_{\mathrm{tail}}$ can be made negligible by choosing $A$ sufficiently large.
Let $\bm{\mathrm{Z}}$ be an $m$-dimensional transverse complete intersection with degree $\leq d$. Let $B_{\rho}$ be a $\rho$-ball with $\rho\gg r^{1/2+{\delta}_{m}}$ and $B_{\rho}\cap N_{r^{1/2+{\delta}_{m}}}\left(\bm{\mathrm{Z}}\right)\cap B\left(0,r\right)\neq\emptyset$. Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\bm{\mathrm{Z}},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B_{\rho}}$, where
${\mathbb{T}}_{B_{\rho}}:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[r\right],T^{l}_{\theta,v}\cap B_{\rho}\neq\emptyset\right\}$
For a vector space $V$ spanned by ${\left\{v_{i}=\left(x_{i,1},x_{i,2},\dots,-x_{i,n}\right)\right\}}^{m}_{i=1}$, we define $V^{+}$ by
$V^{+}:=\left\{\sum^{m}_{i=1}{a_{i}v^{+}_{i}}:a_{i}\in\mathbb{R}\mathrm{,}v^{+}_{i}=\left(x_{i,1},x_{i,2},\dots,x_{i,n}\right),\ \mathrm{f}\mathrm{or\ }1\leq i\leq m\right\}$
We cover $B_{\rho}$ by finitely overlapping balls $B$ of radius $r^{1/2+\delta_{m}}$. For a suitable constant $c>0$, if $cB\cap B_{\rho}\cap\mathbf{Z}\neq\emptyset$, we choose a point $x\in cB\cap B_{\rho}\cap\mathbf{Z}$ and denote $V_{B}:=T_{x}\mathbf{Z}$. Let
${\mathcal{B}}_{a}:=\left\{B:\mathrm{\ }\angle\left(V^{+}_{B},V^{\bot}_{B}\right)\gtrsim K^{-2}\right\}$
${\mathcal{B}}_{b}:=\left\{B:\mathrm{\ }\angle\left(V^{+}_{B},V^{\bot}_{B}\right)\lesssim K^{-2}\right\}$
${\mathbb{T}}_{\bm{\mathrm{Z}},B}:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[\bm{\mathrm{Z}},B\left(0,r\right)\right]:T^{l}_{\theta,v}\cap B\neq\emptyset\right\}$
From the above definition, if $B\in\mathcal{B}_{b}$, then $\angle\!\big(M(\theta),V_{B}^{\perp}\big)$ may be small for some $(\theta,v,l)\in\mathbb{T}_{\mathbf{Z},B}$, and consequently the transverse equidistribution estimates may fail for functions concentrated on $\mathbb{T}_{\mathbf{Z},B}$. Fortunately, there are only few sectors $\theta$ for which $(\theta,v,l)\in\mathbb{T}_{\mathbf{Z},B}$ for some $v,l$.
Lemma 2.17 ([15], Lemma 5.8).
Suppose $B\in{\mathcal{B}}_{b}$. Let
$\mathbb{T}\left[{\theta}^{\prime}\right]:=\left\{\left(\theta,v,l\right):\theta={\theta}^{\prime}\right\},\ \ {\mathrm{\Theta}}_{B}:=\left\{\theta:\mathbb{T}\left[\theta\right]\cap{\mathbb{T}}_{\bm{\mathrm{Z}},B}\neq\emptyset\right\}$
Then ${\mathrm{\Theta}}_{B}$ is contained in the union of $O\left(1\right)$ many $K^{-2}$-sectors.
Now, we recall the definition of the essential part $g_{\mathrm{ess}}$ and the tail part $g_{\mathrm{tail}}$.
${\mathbb{T}}_{\mathrm{ess}}:=\left\{\left({\zeta}_{0},v_{0}\right):\exists\left(\theta,v,l\right)\in{\mathbb{T}}_{{\zeta}_{0},v_{0}}\left.y\right.\mathrm{\ s.t.\ }T^{l}_{\theta,v}\cap B\neq\emptyset\mathrm{\ }\mathrm{fo}\mathrm{r\ some\ }B\in{\mathcal{B}}_{a}\right\}$
${\mathbb{T}}_{\mathrm{tail}}:=\left\{\left({\zeta}_{0},v_{0}\right):\forall\left(\theta,v,l\right)\in{\mathbb{T}}_{{\zeta}_{0},v_{0}}\left.y\right.\mathrm{,\ }T^{l}_{\theta,v}\cap B=\emptyset\mathrm{\ }\mathrm{fo}\mathrm{r\ all\ }B\in{\mathcal{B}}_{a}\right\}$
$g_{\mathrm{ess}}:=\sum_{\left({\zeta}_{0},v_{0}\right)\in{\mathbb{T}}_{\mathrm{ess}}}{g_{{\zeta}_{0},v_{0}}},\ \ g_{\mathrm{tail}}:=\sum_{\left({\zeta}_{0},v_{0}\right)\in{\mathbb{T}}_{\mathrm{tail}}}{g_{{\zeta}_{0},v_{0}}}$ (2.4)
Suppose $A$ is chosen to be sufficiently large. By the definition of $k$-broad norm and Lemma 2.17,
${\left\|Eg_{\mathrm{tail}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}=\mathrm{RapDec}\left(r\right){\left\|g\right\|}_{2}$
Combining this with Lemma 2.11, we have
${\left\|Eg\right\|}_{BL^{p}_{k,A}\left(B_{\rho}\right)}\lesssim{\left\|Eg_{\mathrm{ess}}\right\|}_{BL^{p}_{k,A/2}\left(B_{\rho}\right)}+\mathrm{RapDec}\left(R\right){\left\|g\right\|}_{2}$ (2.5)
As shown in [15], the transverse equidistribution estimates hold for the essential part $g_{\mathrm{ess}}$. The precise statement of the transverse equidistribution estimates for the cone is as follows:
Lemma 2.18 ([15], Lemma 5.13).
Let $B_{\rho}$ be a $\rho$-ball with $\rho\gg r^{1/2+{\delta}_{m}}$ and $B_{\rho}\cap N_{r^{1/2+{\delta}_{m}}}\left(\bm{Z}\right)\cap B\left(0,r\right)\neq\emptyset$. Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\bm{Z},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B_{\rho}}$ and $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$  then
${\left\|g_{\mathrm{ess,}b\mathrm{\ }}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g\right\|}^{2}_{2}$
where $g_{\mathrm{ess,}b\mathrm{\ }}:={\left(g_{\mathrm{ess\ }}\right)}_{b}$.
To apply the nested polynomial Wolff axioms, we need to generalize Lemma 2.18.
Lemma 2.19.
Let $\rho\gg r^{1/2+{\delta}_{m}}$ and $B\left(y,\rho\right)\cap N_{r^{1/2+{\delta}_{m}}}\left(\bm{Z}\right)\cap B\left(0,r\right)\neq\emptyset$. Suppose $g$ is concentrated on the wave packets from $\mathbb{T}\left[\bm{Z},B\left(0,r\right)\right]\cap{\mathbb{T}}_{B\left(y,\rho\right)}$ and $b\in B\left(0,r^{1/2+{\delta}_{m}}\right)$. For $\widetilde{\mathcal{T}}\subseteq\mathbb{T}\left[{\bm{\mathrm{Z}}}_{b}-y,B\left(0,\rho\right)\right]$, there holds
${\left\|{\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g_{\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$
where $\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}$ denotes the set of $\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]$ for which there exists some $\left(\zeta,w,L\right)\in\widetilde{\mathcal{T}}$ satisfying
${\mathrm{D}\mathrm{ist}}_{H}\left(\theta,\ \zeta\right)\leq C{\rho}^{-1/2},\ \ {\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w}\left(y\right),\ T^{l}_{\theta,v}\cap B\left(y,\rho\right)\right)\leq Cr^{1/2+\delta}$
for some constant $C$, and ${\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}:={{\left(g_{\mathrm{ess}}\right)}}_{\widetilde{\mathcal{T}}}^{\widetilde{}}$.
Proof.
Let $h=g_{\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}}$. By Lemma 2.18,
${\left\|{\tilde{h}}_{\mathrm{ess},b}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g_{\mathrm{\uparrow}\mathrm{\uparrow}\widetilde{\mathcal{T}}}\right\|}^{2}_{2}$ (2.6)
By Lemma 2.1 and 2.6,
${\left\|{\tilde{h}}_{\mathrm{ess},b}\right\|}^{2}_{2}\gtrsim{\left\|{\left(h_{\mathrm{ess,\uparrow\tilde{\mathcal{T}}}}\right)}^{\widetilde{}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$ (2.7)
here $\mathrm{\uparrow}\widetilde{\mathcal{T}}$ denotes the set of $\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]$ for which there exists some $\left(\zeta,w,L\right)\in\widetilde{\mathcal{T}}$ satisfying
${\mathrm{D}\mathrm{ist}}_{H}\left(\theta,\ \zeta\right)\leq c{\rho}^{-1/2},\ \ {\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w}\left(y\right),\ T^{l}_{\theta,v}\cap B\left(y,\rho\right)\right)\leq cr^{1/2+\delta}$
for some large constant $c$. Also, Lemma 2.6 gives
${\left\|{\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}\right\|}^{2}_{2}={\left\|{\left(h_{\mathrm{ess,\uparrow\tilde{\mathcal{T}}}}\right)}^{\widetilde{}}_{\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$ (2.8)
Combining (2.6), (2.7), and (2.8), we obtain
${\left\|{\tilde{g}}_{\mathrm{ess,}\widetilde{\mathcal{T}}\mathrm{\ }}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left(r/\rho\right)}^{-\left(n-m\right)/2}{\left\|g_{\mathrm{ess,}\ \mathrm{\uparrow}\widetilde{\mathcal{T}}}\right\|}^{2}_{2}+\mathrm{RapDec}\left(r\right){\left\|g\right\|}^{2}_{2}$
2.8. Polynomial Partitioning
Polynomial partitioning plays a central role in broad estimates. In this subsection, we review the theory of polynomial partitioning. This method is the primary tool used by Guth [5, 6] and Ou–Wang [15] to improve restriction estimates, and it has also led to progress on several related problems, including the Bochner–Riesz conjecture [22] and the Schrödinger maximal estimate [3].
For a polynomial $P:\mathbb{R}^{n}\to\mathbb{R}$, let $\mathrm{cell}(P)$ denote the collection of connected components of $B_{r}\setminus Z(P)$, which partitions the space into finitely many disjoint regions. Each $O\in\mathrm{cell}(P)$ is called a cell cut out by $P$.
Theorem 2.20 ([6], Section 8).
Fix $r\gg 1$, and $d\in{\mathbb{N}}^{+}$. Suppose $F\in L^{1}\left({\mathbb{R}}^{n}\right)$ is supported on $B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\bm{\mathrm{Z}}$, where $\bm{\mathrm{Z}}$ is an $m$-dimensional transverse complete intersection of degree at most $d$. Then, at least one of the following cases holds:
Cellular case. There exists a polynomial $P:{\mathbb{R}}^{n}\to\mathbb{R}$ of degree $O\left(d\right)$ with the following properties:
[…]
(2).For each $O\in\mathcal{O}$, ${\left\|F\right\|}_{L^{1}\left(O\right)}\sim d^{-m}{\left\|F\right\|}_{L^{1}\left(B_{r}\right)}$
where $\mathcal{O}$ is a refinement of the set of connected components of $B_{r}\backslash N_{r^{1/2+{\delta}_{m}}}Z\left(P\right)$.
Algebraic case. There exists a $\left(m-1\right)$-dimensional transverse complete intersection $\bm{\mathrm{Y}}\subseteq\bm{\mathrm{Z}}$ of degree $O\left(d\right)$ such that
${\left\|F\right\|}_{L^{1}\left(B_{r}\right)}\lesssim{\left\|F\right\|}_{L^{1}\left(B_{r}\cap N_{r^{1/2+{\delta}_{m}}}\bm{\mathrm{Y}}\right)}$
We now apply polynomial partitioning to the broad estimates. If the cellular case holds, then
$\|Ef\|^{p}_{BL^{p}_{k,A}\!\left(B_{r}\cap N_{r^{1/2+\delta_{m}}}\mathbf{Z}\right)}\lesssim d^{m}\|Ef\|^{p}_{BL^{p}_{k,A}(O)}\quad\text{for all }O\in\mathcal{O},$
where $\mathcal{O}$ denotes the collection of cells produced by Theorem 2.20. If the algebraic case holds, then
$\|Ef\|^{p}_{BL^{p}_{k,A}\!\left(B_{r}\cap N_{r^{1/2+\delta_{m}}}\mathbf{Z}\right)}\lesssim\|Ef\|^{p}_{BL^{p}_{k,A}\!\left(B_{r}\cap N_{r^{1/2+\delta_{m}}}\mathbf{Y}\right)},$
where $\mathbf{Y}$ is an (m1)(m-1)-dimensional transverse complete intersection produced by Theorem 2.20.
By the fundamental theorem of algebra, we deduce that
Lemma 2.21.
For any $r$-tube $T^{l}_{\theta,v}$,
$\left\{O\in{\mathcal{O}}^{\prime}:O\cap T^{l}_{\theta,v}\neq\emptyset\right\}\lesssim d$
2.9. Nested Polynomial Wolff Axiom.
To refine the work of Ou–Wang [15], we require the nested polynomial Wolff axioms. We begin by recalling the definitions of grains and multigrains from Section 3 of [12].
Definition 2.22 ([12], Definition 3.3).
A grain is defined to be a pair $\left(Z,B\left(y,r\right)\right)$ where $Z$ is a transverse complete intersection and $B_{r}$ is a ball of radius $r>0$. The dimension of a grain $\left(Z,B\left(y,r\right)\right)$ is that of the variety $Z$, while its degree is that of $Z$ and its scale is $r$.
Definition 2.23 ([12], Definition 3.4).
Let $S=\left(Z,B\left(y,r\right)\right)$ be a grain of dimension $m$. A function $f$ with $\mathrm{supp}\ f\subseteq 2{\overline{B}}^{n-1}\backslash B^{n-1}$ is said to be tangent to $S$ if it is concentrated on scale $r$ wave packets belonging to the collection:
$\mathbb{T}\left[S\right]:=\left\{\left(\theta,v,l\right)\in\mathbb{T}\left[r\right]:T^{l}_{\theta,v}\left(y\right)\ \mathrm{is}\ \mathrm{tangent\ to}\ Z\ \mathrm{in}\ B\left(y,r\right)\right\}$
Definition 2.24 ([12], Definition 3.5).
A multigrain is a tuple of grains
$\overrightarrow{S}=\left(S_{s},S_{s+1},\dots,S_{t}\right),\ \ S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$
satisfying
${\mathrm{dim}Z_{i}\ }=i,\quad\forall i$
$B(y_{i},r_{i})\cap B(y_{i+1},r_{i+1})\neq\emptyset,\quad\forall i$
$N_{Cr^{1/2+{\delta}_{s}}_{s}}Z_{s}\subseteq\dots\subseteq N_{Cr^{1/2+{\delta}_{t}}_{t}}Z_{t}\mathrm{\ for\ some\ }C>0$
$0<r_{s}\leq r_{s+1}\leq\dots\leq r_{t}$
The parameter (ts+1)(t-s+1) is called the multigrain level. The degree of $\overrightarrow{S}$ is defined as $\max_{1\leq i\leq l}Z_{i}$. The multiscale of $\overrightarrow{S}$ is the tuple $\overrightarrow{r}=(r_{s},r_{s+1},\dots,r_{t})$. Finally, we write
${\overrightarrow{S}}_{1}=\left(S_{s},S_{s+1},\dots,S_{t}\right)\preccurlyeq{\overrightarrow{S}}_{2}=\left(S_{s^{\prime}},S_{s^{\prime}+1},\dots,S_{t}\right),\ \ \mathrm{if\ }s^{\prime}\geq s$
The restriction conjecture is closely related to a Kakeya-type problem concerning the continuum incidence theory of tubes. More precisely, let $\overrightarrow{S}_{s,t}=(S_{s},\dots,S_{t})$ be a multigrain with $S_{i}=(Z_{i},B(y_{i},r_{i}))$. Suppose $\mathcal{L}$ is a collection of $r_{t}^{-1/2}$-direction-separated rays $L$ satisfying
$L\cap B(y_{i},r_{i})\cap N_{\rho_{j}}Z_{i}\neq\emptyset$
and
$L\cap B(y_{i},2r_{i})\subseteq N_{\rho_{j}}Z_{i}\cap B(y_{i},2r_{i})$
then what is the maximal possible cardinality of $\mathcal{L}$? Understanding this incidence problem is crucial for obtaining effective estimates for $Ef$. It is well known that the restriction conjecture implies the Kakeya conjecture, the latter providing a heuristic bound on the number of incidences between direction-separated tubes.
In [12, 8, 10], a multiscale structure, referred to as nested tubes, is introduced to describe the spatial relationships between tubes at different scales.
Definition 2.25 (Nested Tubes).
Let ${\overrightarrow{S}}_{m}=\left(S_{m},\dots,S_{n}\right)$ be a multigrain with $S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$ and ${\mathrm{dim}Z_{i}\ }=i$. We define $\mathbb{T}\left[{\overrightarrow{S}}^{i}_{m}\right]$ by induction on $i\in\mathbb{N}\cap\left[m,n-1\right]$. For $i=m$, we set
$\mathbb{T}\left[{\overrightarrow{S}}^{m}_{m}\right]:=\mathbb{T}\left[S_{m}\right]\ \ \mathrm{\Theta}\left[{\overrightarrow{S}}^{m}_{m}\right]:=\left\{\theta:\exists T^{l}_{\theta,v}\in\mathbb{T}\left[{\overrightarrow{S}}^{m}_{m}\right]\ \mathrm{for\ some\ }v,l\right\}$
For $m<i\leq n$, let $\mathbb{T}\left[{\overrightarrow{S}}^{i}_{m}\right]$ be the set of all $r_{i}$-tubes $T^{l}_{\theta,v}$ satisfying the following conditions:
(a). $T^{l}_{\theta,v}\in\mathbb{T}\left[Z_{i},B\left(y_{i},r_{i}\right)\right]\$
(b). There exists $\exists T^{L}_{\zeta,w}\in\mathbb{T}\left[Z_{i-1},B\left(y_{i-1},r_{i-1}\right)\right]$ such that
$\quad{\mathrm{D}\mathrm{ist}}_{H}\left(w,\ \zeta\right)\lesssim r^{-1/2}_{i-1},\quad{\mathrm{D}\mathrm{ist}}_{H}\left(T^{L}_{\zeta,w},\ T^{l}_{\theta,v}\cap B\left(y_{i},r_{i}\right)\right)\lesssim r^{1/2+\delta}_{i}$
and
[…]
Finally, we set
$\mathbb{T}\left[{\overrightarrow{S}}^{\#}_{m}\right]:=\mathbb{T}\left[{\overrightarrow{S}}^{n-1}_{m}\right],\ \ \mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{m}\right]:=\ \mathrm{\Theta}\left[{\overrightarrow{S}}^{n-1}_{m}\right]$
In [12, 8, 10], a result on nested tubes as a corollary of Lemma 2.11 in [25]—was used to obtain new results on the Bochner–Riesz problem, Stein’s square function, and Fourier restriction for the paraboloid. We state this result below.
Lemma 2.26 ([12], Lemma 3.7, Nested polynomial Wolff axioms).
Let ${\overrightarrow{S}}_{m}=\left(S_{m},\dots,S_{n}\right)$ be a multigrain with $S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$ such that for all $i\in\mathbb{N}\cap\left[m,n\right]$,
${\mathrm{dim}Z_{i}\ }=i\ ,\ \ {\mathrm{deg}Z_{i}\ }\leq d$
then
$\mathrm{\Theta}\left[{\overrightarrow{S}}^{n}_{m}\right]{\lesssim}_{d}r^{O\left({\delta}_{0}\right)}_{n}\left(\prod^{n-1}_{j=m}{r^{-\frac{1}{2}}_{j}}\right)r^{\frac{n-1}{2}}_{n}$
Consequently,
Corollary 2.27.
Let ${\overrightarrow{S}}_{m}=\left(S_{m},\dots,S_{n}\right)$ be a multigrain with $S_{i}=\left(Z_{i},B\left(y_{i},r_{i}\right)\right)$ such that for all $i\in\mathbb{N}\cap\left[m,n\right]$,
${\mathrm{dim}Z_{i}\ }=i\ ,\ \ {\mathrm{deg}Z_{i}\ }\leq d$
then
$\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{m}\right]{\lesssim}_{d}r^{O\left({\delta}_{0}\right)}_{n-1}\left(\prod^{n-2}_{j=m}{r^{-\frac{1}{2}}_{j}}\right)r^{\frac{n-2}{2}}_{n-1}$
Proof.
Apply Lemma 2.26 to ${\overrightarrow{S}}:=(S_{m},...,S_{n-1},({\mathbb{R}}^{n},B(y_{n-1},r_{n-1})))$. ∎
3. Finding Polynomial Structure
In this section, we reformulate the induction arguments of [15] as a recursive procedure, combining ideas from [11, 12]. The argument consists of two algorithms.
In Algorithm 1, we iteratively apply polynomial partitioning until the tangential term dominates. This procedure reduces the problem from an $m$-dimensional to an (m1)(m-1)-dimensional setting and corresponds to the induction on scale. Algorithm 2 is obtained by repeatedly applying Algorithm 1, and the iteration terminates once the minimal dimension is reached.
Throughout the iteration, we assume without loss of generality that the error terms $\mathrm{RapDec}(R)\|f\|_{2}$ are negligible.
3.1. Algorithm 1
This algorithm proceeds via iterated polynomial partitioning. At each step, one of three cases occurs: the cellular case, the transverse case, or the tangential case. The iteration terminates either when the tangential case occurs or when the scale becomes sufficiently small.
Input. Algorithm 1 takes the following data as input:
  • A grain $(\mathbf{Z},B(y,r))$ of dimension $m$.
  • A function $f\in L^{\infty}\!\left(2\overline{B}^{\,n-1}\setminus B^{n-1}\right)$ that is tangent to $(\mathbf{Z},B(y,r))$.
  • An admissible integer $A\in\mathbb{N}$.
At the initial step, we set
$O_{0}:=\{\,B_{r}\cap N_{r^{1/2+\delta_{m}}}\mathbf{Z}\,\},\qquad f_{O_{0}}:=f,\qquad\mathfrak{h}_{0}:=\emptyset,\qquad\rho_{0}:=r.$
Output. At the $j$-th iteration, Algorithm 1 outputs:
  • History. A word $\mathfrak{h}_{j}$ of length $j$ in the alphabet $\{\mathrm{a},\mathrm{c}\}$, recording how each cell $O_{j}\in\mathcal{O}_{j}$ is produced by the iterated polynomial partitioning. Let $\mathrm{a}(j)$ and $\mathrm{c}(j)$ denote the number of letters $\mathrm{a}$ and $\mathrm{c}$ in $\mathfrak{h}_{j}$, respectively.
  • Spatial scale $\rho_{j}$. The scale $\rho_{j}$ is determined by the initial scale $r$ and the history $\mathfrak{h}_{j}$. Define an auxiliary parameter $\widetilde{\delta}_{m-1}$ by
    $(1-\widetilde{\delta}_{m-1})\!\left(\frac{1}{2}+\delta_{m-1}\right)=\frac{1}{2}+\delta_{m},$
    which implies
    $\frac{\delta_{m-1}}{2}\leq\widetilde{\delta}_{m-1}\leq 2\delta_{m-1}.$
  • Cells. A family of subsets $\mathcal{O}_{j}\subset\mathbb{R}^{n}$. Each cell $O_{j}\in\mathcal{O}_{j}$ is contained in a ball $B_{O_{j}}=B(y_{O_{j}},\rho_{j})$.
  • Localized functions. A collection of functions $\{f_{O_{j}}\}_{O_{j}\in\mathcal{O}_{j}}$. For each cell $O_{j}$ there exists a translate $\mathbf{Z}_{O_{j}}:=\mathbf{Z}+x_{O_{j}}$ such that $f_{O_{j}}$ is tangent to the grain $(\mathbf{Z}_{O_{j}},B_{O_{j}})$.
  • Auxiliary parameters. A large integer $d$ depending only on the admissible parameters and $\deg\mathbf{Z}$, together with
    $C^{\mathrm{I}}_{j,\delta}(d,r):=d^{\#\mathrm{c}(j)\delta}\,(\log r)^{\#\mathrm{a}(j)(1+\delta)},\qquad C^{\mathrm{II}}_{j,\delta}(d):=d^{\#\mathrm{c}(j)\delta+n\#\mathrm{a}(j)(1+\delta)},$
    $C^{\mathrm{III}}_{j,\delta}(d,r):=d^{j\delta}r^{\overline{C}\#\mathrm{a}(j)\delta_{m}},\qquad C^{\mathrm{IV}}_{j,\delta}(d,r):=C^{\mathrm{III}}_{j,\delta}(d,r),\qquad A_{j}:=3^{-\#\mathrm{a}(j)}A.$
Properties. At the $j$-th iteration:
  1. (1)
    The mass of $\|Ef\|^{p}_{BL^{p}_{k,A}(B_{r})}$ is concentrated on the cells $O_{j}\in\mathcal{O}_{j}$:
    $\|Ef\|^{p}_{BL^{p}_{k,A}(B_{r})}\leq C^{\mathrm{I}}_{j,\delta}(d,r)\sum_{O_{j}\in\mathcal{O}_{j}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A}(O_{j})}.$ (3.1)
  2. (2)
    The functions $f_{O_{j}}$ satisfy
    $\sum_{O_{j}\in\mathcal{O}_{j}}\|f_{O_{j}}\|_{2}^{2}\leq C^{\mathrm{II}}_{j,\delta}(d)\,d^{\#\mathrm{c}(j)}\|f\|_{2}^{2}.$ (3.2)
  3. (3)
    Each $f_{O_{j}}$ satisfies, for all $1\leq\rho\leq\rho_{j}$,
    $\|f_{O_{j}}\|_{2}^{2}\leq C^{\mathrm{III}}_{j,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}d^{-\#\mathrm{c}(j)(m-1)}\|f\|_{2}^{2},$ (3.3)
    and for all $\rho\in[1,\rho_{j}]$,
    $\max_{\theta:\,\rho^{-1/2}\text{-sector}}\|f_{O_{j}}\|_{L^{2}_{\mathrm{avg}}(\theta)}^{2}\leq\\ C^{\mathrm{III}}_{j,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}\max_{\theta:\,\rho^{-1/2}\text{-sector}}\|f\|_{L^{2}_{\mathrm{avg}}(\theta)}^{2}$ (3.4)
  4. (4)
    For $\mathcal{T}\subseteq\mathbb{T}[\rho_{j}]$ and $O_{j}\in\mathcal{O}_{j}$,
    $\|f_{O_{j},\mathcal{T}}\|_{2}^{2}\leq C^{\mathrm{IV}}_{j,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}\|f_{\uparrow^{j}\mathcal{T}}\|_{2}^{2},$ (3.5)
    where $(f_{O_{j},\mathcal{T}})^{\widetilde{}}:=({f_{O_{j}}})^{\widetilde{}}_{\mathcal{T}}$, and $\uparrow^{j}\mathcal{T}$ denotes the set of wave packets $(\theta,v,l)\in\mathbb{T}[r]$ such that
    $\mathrm{dist}_{H}(\theta,\zeta)\leq c_{j}\rho_{j}^{-1/2}$
    $\mathrm{dist}_{H}\!\left(T_{\zeta,w}^{L}(y_{O_{j}}),\,T_{\theta,v}^{l}(y)\cap B_{O_{j}}\right)\leq c_{j}r^{1/2+\delta}$
    for some $(\zeta,w,L)\in\mathcal{T}$, where
    $c_{j}:=\sum_{i=1}^{j}2^{-(j-1)/2}C$
    for an absolute constant $C$.
Recursion Steps. If $\rho_{j}\leq r^{\widetilde{\delta}_{m-1}}$, then the algorithm terminates. In this case we set $J:=j$ and $\mathcal{O}:=\mathcal{O}_{j}$, and label this outcome as [tiny]. Otherwise, for each $O_{j}\in\mathcal{O}_{j}$ we apply polynomial partitioning with degree $d$ to $\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A}(O_{j})}$.
Let $\mathcal{O}_{j,\mathrm{cell}}\subseteq\mathcal{O}_{j}$ denote the collection of cells for which the cellular case occurs, and set
$\mathcal{O}_{j,\mathrm{alg}}:=\mathcal{O}_{j}\setminus\mathcal{O}_{j,\mathrm{cell}}.$
then one of the following cases holds:
Cellular-dominant case. Suppose
$\sum_{O_{j}\in\mathcal{O}_{j,\mathrm{alg}}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A}(O_{j})}\leq\sum_{O_{j}\in\mathcal{O}_{j,\mathrm{cell}}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A}(O_{j})}.$ (3.6)
Append the letter $\mathrm{c}$ to $\mathfrak{h}_{j}$ to obtain $\mathfrak{h}_{j+1}$. Then
$\mathrm{a}(j+1)=\#\mathrm{a}(j),\qquad\#\mathrm{c}(j+1)=\#\mathrm{c}(j)+1,\qquad A_{j+1}=A_{j},$
and
$C^{\mathrm{I}}_{j+1,\delta}=d^{\delta}C^{\mathrm{I}}_{j,\delta},\qquad C^{\mathrm{II}}_{j+1,\delta}=d^{\delta}C^{\mathrm{II}}_{j,\delta},$
$C^{\mathrm{III}}_{j+1,\delta}=d^{\delta}C^{\mathrm{III}}_{j,\delta},\qquad C^{\mathrm{IV}}_{j+1,\delta}=d^{\delta}C^{\mathrm{IV}}_{j,\delta}.$
Set $\rho_{j+1}:=\rho_{j}/2$. For each $O_{j}\in\mathcal{O}_{j,\mathrm{cell}}$, the cellular case yields a polynomial $P$ of degree $O(d)$ such that:
  1. (1)
    $\mathcal{O}(O_{j})\sim d^{m}$,
  2. (2)
    every $O\in\mathcal{O}(O_{j})$ lies in a ball $B_{O}=B(y_{O},\rho_{j+1})$,
  3. (3)
    $\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A_{j}}(O)}\sim d^{-m}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A_{j}}(O_{j})}$.
Here $\mathcal{O}(O_{j})$ refines the connected components of $O_{j}\setminus N_{\rho_{j}^{1/2+\delta_{m}}}Z(P)$. Define
$f_{O}:=\sum_{\begin{subarray}{c}(\theta,v,l)\in\mathbb{T}[\rho_{j}]\\ T_{\theta,v}^{l}\cap O\neq\emptyset\end{subarray}}(f_{O_{j}})_{\theta,v}^{l}.$
By Lemma 2.21 and Lemma 2.1,
$\sum_{O\in\mathcal{O}(O_{j})}\|f_{O}\|_{2}^{2}\lesssim d\,\|f_{O_{j}}\|_{2}^{2}.$
Pigeonholing gives a refinement $\mathcal{O}^{\prime}(O_{j})$ with $\mathcal{O}^{\prime}\sim d^{m}$ such that
$\|f_{O}\|_{2}^{2}\lesssim d^{-m+1}\|f_{O_{j}}\|_{2}^{2}.$
Since $\rho_{j+1}\sim\rho_{j}$, there exists a finite set $\mathcal{B}\subset B(0,\rho_{j}^{1/2+\delta_{m}})$ with $\mathcal{B}\sim 1$ such that
$N_{\rho_{j}^{1/2+\delta_{m}}}\mathbf{Z}_{O_{j}}\subset\bigcup_{b\in\mathcal{B}}N_{\rho_{j+1}^{1/2+\delta_{m}}}(\mathbf{Z}_{O_{j}}+b).$
For $O\in\mathcal{O}^{\prime}(O_{j})$ and $b\in\mathcal{B}$ define
$O_{b}:=O\cap N_{\rho_{j+1}^{1/2+\delta_{m}}}(\mathbf{Z}_{O_{j}}+b),$
$\widetilde{f_{O_{b}}}:=\sum_{\begin{subarray}{c}(\zeta,w,L)\in\mathbb{T}[\rho_{j+1}]\\ T_{\zeta,w}^{L}(y_{O})\cap O_{b}\neq\emptyset\end{subarray}}(\widetilde{f_{O}})_{\zeta,w}^{L},\qquad\widetilde{f_{O}}(\xi)=e^{i\psi_{y_{O}}(\xi)}f_{O}(\xi).$
By Lemma 2.6, $f_{O_{b}}$ is tangent to $(\mathbf{Z}_{O_{j}}+b,B_{O})$. Set
$\mathcal{O}_{j+1}:=\bigcup_{O_{j}\in\mathcal{O}_{j,\mathrm{cell}}}\bigcup_{O\in\mathcal{O}^{\prime}(O_{j})}\{O_{b}:\,b\in\mathcal{B}\}.$
Properties 1–3 follow from [11]. We verify Property 4. By Lemma 2.1 and Lemma 2.6, for $\mathcal{T}\subseteq\mathbb{T}[\rho_{j+1}]$,
$\|\widetilde{f}_{O_{j+1},\mathcal{T}}\|_{2}^{2}\lesssim\|f_{O_{j},\uparrow\mathcal{T}}\|_{2}^{2}.$ (3.7)
Here $\uparrow\mathcal{T}$ denotes the set of wave packets $(\theta,v,l)\in\mathbb{T}[\rho_{j}]$ such that
$\mathrm{Dist}_{H}(\theta,\zeta)\leq C\rho_{j+1}^{-1/2},\qquad\mathrm{Dist}_{H}\!\left(T_{\zeta,w}^{L}(y_{O_{j+1}}),\,T_{\theta,v}^{l}\cap B(y_{O_{j+1}},\rho_{j+1})\right)\leq C\rho_{j}^{1/2+\delta}$
for some $(\zeta,w,L)\in\mathcal{T}$, where $C$ is the absolute constant from Property 4.Combining (3.5) and (3.7), we obtain
$\|f_{O_{j+1},\mathcal{T}}\|_{2}^{2}\lesssim d^{-\delta}C^{\mathrm{IV}}_{j+1,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}\|f_{\uparrow^{j}(\uparrow\mathcal{T})}\|_{2}^{2}.$
Since $\rho_{j+1}\sim\rho_{j}$, choosing $d$ sufficiently large yields
$\|f_{O_{j+1},\mathcal{T}}\|_{2}^{2}\leq C^{\mathrm{IV}}_{j+1,\delta}(d,r)\left(\frac{r}{\rho_{j+1}}\right)^{-(n-m)/2}\|f_{\uparrow^{j}(\uparrow\mathcal{T})}\|_{2}^{2}.$
From [12] we see ${\uparrow}^{j}\left(\uparrow\mathcal{T}\right)\subseteq{\uparrow}^{j+1}\mathcal{T}$. This completes the verification.
Algebraic-dominant Case. If (3.6) fails, then
$\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,cell}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}\leq\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A}\left(O_{j}\right)}}$
Append the letter $\mathrm{a}$ to $\mathfrak{h}_{j}$ to obtain $\mathfrak{h}_{j+1}$. Now we have
$\mathrm{a}\left(j+1\right)=\#\mathrm{a}\left(j\right)+1,\ \ \#\mathrm{c}\left(j+1\right)=\#\mathrm{c}\left(j\right),\ \ A_{j+1}=A_{j}/3$
$C^{\mathrm{I}}_{j+1,\delta}\left(d,r\right)={{\mathrm{log}}^{1+\delta}r\ }C^{\mathrm{I}}_{j,\delta}\left(d,r\right),\ \ C^{\mathrm{II}}_{j+1,\delta}\left(d\right)=d^{n\left(1+\delta\right)}C^{\mathrm{II}}_{j,\delta}\left(d\right)$
$C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right)=r^{\overline{C}{\delta}_{m}}d^{\delta}C^{\mathrm{III}}_{j,\delta}\left(d,r\right),\ \ C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right)=r^{\overline{C}{\delta}_{m}}d^{\delta}C^{\mathrm{IV}}_{j,\delta}\left(d,r\right)$
Set $\rho_{j+1}:=\rho_{j}^{\,1-\widetilde{\delta}_{m-1}}$. Since the algebraic case occurs, for each $O_{j}\in\mathcal{O}_{j,\mathrm{alg}}$ there exists an (m1)(m-1)-dimensional transverse complete intersection $\mathbf{Y}\subseteq\mathbf{Z}_{O_{j}}$ of degree $C_{\mathrm{alg}}d$ such that
${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}\lesssim{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\bm{\mathrm{Y}}\right)}$
where $C_{\mathrm{alg}}$ is a sufficiently large constant. Let $\mathcal{B}_{O_{j}}$ denote the collection of such balls. For each $B\in\mathcal{B}_{O_{j}}$, define
$\mathbb{T}_{B}:=\left\{(\theta,v,l)\in\mathbb{T}[\rho_{j}]:T_{\theta,v}^{l}\cap B\cap N_{\rho_{j}^{1/2+\delta_{m}}}\mathbf{Y}\neq\emptyset\right\},\qquad f_{B}:=(f_{O_{j}})_{\mathbb{T}_{B}}.$
Let $\mathbb{T}_{B,\mathrm{tang}}\subseteq\mathbb{T}_{B}$ be the set of $\rho_{j}$-scale wave packets $(\theta,v,l)$ satisfying:
  1. (1)
    $T_{\theta,v}^{l}\cap 2B\subseteq N_{\rho_{j}^{1/2+\delta_{m}}}\mathbf{Y}=N_{\rho_{j+1}^{1/2+\delta_{m-1}}}\mathbf{Y}$;
  2. (2)
    whenever $x\in T_{\theta,v}^{l}$ and $y\in\mathbf{Y}\cap 2B$ satisfy $|x-y|\lesssim\rho_{j}^{1/2+\delta_{m}}=\rho_{j+1}^{1/2+\delta_{m-1}}$, then
    $\angle\!\big(L(\theta),T_{y}\mathbf{Y}\big)\lesssim\rho_{j+1}^{-1/2+\delta_{m-1}}.$
Define $\mathbb{T}_{B,\mathrm{trans}}:=\mathbb{T}_{B}\setminus\mathbb{T}_{B,\mathrm{tang}}$ and
$f_{B,\mathrm{tang}}:=(f_{O_{j}})_{\mathbb{T}_{B,\mathrm{tang}}},\qquad f_{B,\mathrm{trans}}:=(f_{O_{j}})_{\mathbb{T}_{B,\mathrm{trans}}}.$
By Lemma 2.12,
${\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}\ }\left(O_{j}\right)}\lesssim\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{tang}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}}+\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{trans}}\right\|}^{p}_{BL^{p}_{k,2A_{j+1}}\left(B\right)}}$ (3.8)
By a pigeonholing argument, at least one of the following cases holds.
Tangential Subcase. Let $C_{\mathrm{tang}}$ be a large constant. If
$\sum_{O_{j}\in\mathcal{O}_{j,\mathrm{alg}}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A_{j}}(O_{j})}\leq\frac{C_{\mathrm{tang}}}{2}\sum_{O_{j}\in\mathcal{O}_{j,\mathrm{alg}}}\sum_{B\in\mathcal{B}_{O_{j}}}\|Ef_{B,\mathrm{tang}}\|^{p}_{BL^{p}_{k,A_{j+1}}(B)},$ (3.9)
then the algorithm terminates. We set $J=j+1$ and label this outcome [tang]. Define the collection of grains
$\mathcal{S}:=\left\{(\mathbf{Y},B):B\in\mathcal{B}_{O_{j}}\ \text{for some}\ O_{j}\in\mathcal{O}_{j,\mathrm{alg}}\right\}.$
For each $S=(\mathbf{Y},B)\in\mathcal{S}$, set $f_{S}:=f_{B,\mathrm{tang}}$ and $B_{S}:=B$. Lemma 2.6 implies that $f_{S}$ is tangent to $S$. Since (3.6) fails, inequality (3.9) yields
$\sum_{O_{j}\in\mathcal{O}_{j}}\|Ef_{O_{j}}\|^{p}_{BL^{p}_{k,A_{j}}(O_{j})}\leq C_{\mathrm{tang}}\sum_{S\in\mathcal{S}}\|Ef_{S}\|^{p}_{BL^{p}_{k,A_{j+1}}(B_{S})}.$ (3.10)
By Lemma 2.1 and Corollary 2.3, for all $1\leq\rho\leq\rho_{j+1}$, we have
${\mathop{\mathrm{max}}_{S\in\mathcal{S}}{\left\|f_{S}\right\|}^{2}_{2}\ }\leq C_{\mathrm{tang}}\mathop{\mathrm{max}}_{O_{j}\in{\mathcal{O}}_{j}}{\left\|f_{O_{j}}\right\|}^{2}_{2}$ (3.11)
${\mathop{\mathrm{max}}_{\begin{array}[]{c}S\in\mathcal{S}\\ \theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}\end{array}}{\left\|f_{S}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\leq C_{\mathrm{tang}}{\mathop{\mathrm{max}}_{\begin{array}[]{c}O_{j}\in{\mathcal{O}}_{j}\\ \theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}\end{array}}{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (3.12)
Since ${\mathcal{B}}_{O_{j}}\lesssim r^{n{\widetilde{\delta}}_{m-1}}$ for all $O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}$, Lemma 2.1 gives
$\sum_{S\in\mathcal{S}}{{\left\|f_{S}\right\|}^{2}_{2}}\leq C_{\mathrm{tang}}r^{n{\widetilde{\delta}}_{m-1}}\sum_{O_{j}\in{\mathcal{O}}_{j}}{{\left\|f_{O_{j}}\right\|}^{2}_{2}}$ (3.13)
Moreover, for $\mathcal{T}\subseteq\mathbb{T}[\rho_{j+1}]$ and $S=(\mathbf{Y},B)\in\mathcal{S}$ with $B\in\mathcal{B}_{O_{j}}$, Lemma 2.1 together with Corollary 2.7 yields
${\left\|f_{S,\mathcal{T}}\right\|}^{2}_{2}\leq C_{\mathrm{tang}}{\left\|f_{O_{j},\uparrow\mathcal{T}}\right\|}^{2}_{2}$ (3.14)
where $\big(f_{S,\mathcal{T}}\big)^{\widetilde{}}:={\left(f_{S}\right)}_{\mathcal{T}}^{\widetilde{}}$
Transverse Subcase: If (3.9) fails, then (3.8) gives
$\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}}\lesssim\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{t}\mathrm{rans}}\right\|}^{p}_{BL^{p}_{k,2A_{j+1}}\left(B\right)}}}$
Let $f_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}:={\left(f_{B,\mathrm{t}\mathrm{rans}}\right)}_{\mathrm{ess}}$. By (2.5),
$\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{{\left\|Ef_{O_{j}}\right\|}^{p}_{BL^{p}_{k,A_{j}}\left(O_{j}\right)}}\lesssim\sum_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{\sum_{B\in{\mathcal{B}}_{O_{j}}}{{\left\|Ef_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}}}$
Fix $O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}$ and $B\in{\mathcal{B}}_{O_{j}}$. Arguing as in [6], for each $b\in{\mathfrak{b}}_{B,O_{j}}$ there exist ${\mathfrak{b}}_{B,O_{j}}\subseteq B\left(0,{\rho}^{1/2+{\delta}_{m}}_{j}\right)$ and $g_{b}$ tangent to $\left({\bm{\mathrm{Z}}}_{O_{j}}+b,B\right)$ such that
${\left\|Ef_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\right)}\lesssim{\mathrm{log}r\ }\sum_{b\in{\mathfrak{b}}_{B,O_{j}}}{{\left\|Eg_{b}\right\|}^{p}_{BL^{p}_{k,A_{j+1}}\left(B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j}}\left({\bm{\mathrm{Z}}}_{O_{j}}+b\right)\right)}}$
$\sum_{b\in{\mathfrak{b}}_{B,O_{j}}}{{\left\|g_{b}\right\|}^{2}_{2}}\lesssim{\left\|f_{B,\mathrm{t}\mathrm{rans}\mathrm{,ess}}\right\|}^{2}_{2}$
Let
$\mathcal{O}\left(O_{j}\right):=\left\{B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\bm{\mathrm{Z}}}_{O_{j}}+b\right):b\in{\mathfrak{b}}_{B,O_{j}},\ B\in{\mathcal{B}}_{O_{j}}\right\}$
For $O=B\cap N_{{\rho}^{1/2+{\delta}_{m}}_{j+1}}\left({\bm{\mathrm{Z}}}_{O_{j}}+b\right)\in\mathcal{O}\left(O_{j}\right)$, define $f_{O}:=g_{b}$. Finally, set
${\mathcal{O}}_{j+1}:=\bigcup_{O_{j}\in{\mathcal{O}}_{j\mathrm{,alg}}}{\mathcal{O}\left(O_{j}\right)}$
We now verify the properties. Properties 1 and 2 follow from [11]; here we verify Properties 3 and 4.
Property 3. By Lemma 2.18,
${\left\|f_{O_{j+1}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j+1}/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j}}\right\|}^{2}_{2}$
Combining this with (3.3), we obtain
$\|f_{O_{j+1}}\|_{2}^{2}\lesssim r^{O(\delta_{m})-\overline{C}\delta_{m}}d^{-\delta}C^{\mathrm{III}}_{j+1,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}d^{-\#\mathrm{c}(j+1)(m-1)}\|f\|_{2}^{2}.$
The induction closes provided $\overline{C}$ and $d$ are chosen sufficiently large. For a $\rho^{-1/2}$-sector $\theta$ with $1\leq\rho\leq\rho_{j+1}$, Lemma 2.6, Lemma 2.19, and Corollary 2.2 yield
${\left\|f_{O_{j+1}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j+1}/{\rho}_{j}\right)}^{-\left(n-m\right)/2}\sum_{\begin{array}[]{c}{\theta}^{\prime}:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}\\ {\theta}^{\prime}\cap C\theta\neq\emptyset\end{array}}{{\left\|f_{O_{j}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left({\theta}^{\prime}\right)}}$
By pigeonholing, there exists a $\rho^{-1/2}$-sector $\theta^{\prime}$ such that $\theta^{\prime}\cap C\theta\neq\emptyset$ and
$\|f_{O_{j+1}}\|^{2}_{L^{2}_{\mathrm{avg}}(\theta)}\lesssim r^{O(\delta_{m})}\left(\frac{\rho_{j+1}}{\rho_{j}}\right)^{-(n-m)/2}\|f_{O_{j}}\|^{2}_{L^{2}_{\mathrm{avg}}(\theta^{\prime})}$
Combining this and (3.4), we obtain
$\max_{\theta:\,\rho^{-1/2}\text{-sector}}\|f_{O_{j+1}}\|^{2}_{L^{2}_{\mathrm{avg}}(\theta)}\lesssim\\ r^{O(\delta_{m})-\overline{C}\delta_{m}}d^{-\delta}C^{\mathrm{III}}_{j+1,\delta}(d,r)\left(\frac{r}{\rho_{j}}\right)^{-(n-m)/2}\max_{\theta:\,\rho^{-1/2}\text{-sector}}\|f\|^{2}_{L^{2}_{\mathrm{avg}}(\theta)}$
Since $\overline{C}$ and $d$ are chosen sufficiently large, this verifies Property 3.
Property 4. For $\mathcal{T}\subseteq\mathbb{T}\left[{\rho}_{j+1}\right]$, Lemma 2.19 implies
${\left\|{\tilde{f}}_{O_{j+1,\mathcal{T}}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)}{\left({\rho}_{j}/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}{\left\|f_{O_{j},\uparrow\mathcal{T}}\right\|}^{2}_{2}$
Combining this with (3.5) we have
${\left\|f_{O_{j+1,\mathcal{T}}}\right\|}^{2}_{2}\lesssim r^{O\left({\delta}_{m}\right)-\overline{C}{\delta}_{m}}d^{-\delta}{\left({\rho}_{j}/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}C^{\mathrm{IV}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\left(\uparrow\mathcal{T}\right)}\right\|}^{2}_{2}$
Since $\overline{C}$ and $d$ are chosen sufficiently large, one has
${\left\|f_{O_{j+1},\ \mathcal{T}}\right\|}^{2}_{2}\leq C^{\mathrm{III}}_{j+1,\delta}\left(d,r\right){\left(r/{\rho}_{j+1}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{j}\left(\uparrow\mathcal{T}\right)}\right\|}^{2}_{2}$
As in the cellular case, we have $\uparrow^{j}(\uparrow\mathcal{T})\subseteq\uparrow^{j+1}\mathcal{T}$, and thus the induction closes.
Summary. We now summarize the output in a concise form. Since
$\mathrm{c}\left(J\right)\lesssim\log{r},\ \ \#\mathrm{a}\left(J\right)\lesssim{\delta}^{-1}_{m-1}\log{{\delta}^{-1}_{m-1}}$ (3.15)
the accumulated error remains negligible after the algorithm terminates. Let $D:=d^{c\left(J\right)}$. If [tang] happens, then (3.1), (3.2), (3.3), (3.4), (3.5), (3.10), (3.11), (3.12), (3.13), and (3.14) gives
${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\lessapprox\sum_{S\in\mathcal{S}}{{\left\|Ef_{S}\right\|}^{p}_{BL^{p}_{k,A_{J}}\left(B_{S}\right)}}$ (3.16)
$\sum_{S\in\mathcal{S}}{{\left\|f_{S}\right\|}^{2}_{2}}\lessapprox D{\left\|f\right\|}^{2}_{2}$ (3.17)
${\mathop{\mathrm{max}}_{S\in\mathcal{S}}{\left\|f_{S}\right\|}^{2}_{2}\ }\lessapprox D^{-(m-1)}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f\right\|}^{2}_{2}$ (3.18)
${\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f_{S}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\mathop{\mathrm{max}}_{\theta:{\rho}^{-1/2}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (3.19)
${\left\|f_{S,\ \mathcal{T}}\right\|}^{2}_{2}\lessapprox{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f_{{\uparrow}^{J}\mathcal{T}}\right\|}^{2}_{2}$ (3.20)
If [tiny] happens, then (3.1), (3.2), (3.3), (3.4), and (3.5) gives
${\left\|Ef\right\|}^{p}_{BL^{p}_{k,A}\left(B_{r}\right)}\lessapprox\sum_{O\in\mathcal{O}}{{\left\|Ef_{O}\right\|}^{p}_{BL^{p}_{k,A_{J}}\left(O\right)}}$ (3.21)
$\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\lessapprox D{\left\|f\right\|}^{2}_{2}$ (3.22)
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox D^{-(m-1)}{\left(r/{\rho}_{J}\right)}^{-\left(n-m\right)/2}{\left\|f\right\|}^{2}_{2}$ (3.23)
3.2. Algorithm 2
Algorithm 2 is obtained by repeatedly applying Algorithm 1 and terminates when the [tiny] case dominates. Consider a family of Lebesgue exponents $p_{i}$ for $k\leq l\leq n$, which will be determined in Section 4, satisfying
$p_{k}\geq p_{k+1}\geq\dots\geq p_{n}=:p>2$ (3.24)
and define $0\leq{\alpha}_{i},{\beta}_{i}\leq 1$ by ${\alpha}_{n}={\beta}_{n}:=1$ and
${\alpha}_{i}:=\left(\frac{1}{2}-\frac{1}{p_{i+1}}\right){\left(\frac{1}{2}-\frac{1}{p_{i}}\right)}^{-1},\ \ {\beta}_{i}:=\left(\frac{1}{2}-\frac{1}{p}\right){\left(\frac{1}{2}-\frac{1}{p_{i}}\right)}^{-1}\ \mathrm{for}\ k\leq i\leq n-1$ (3.25)
Input: A function $f:2{\overline{B}}^{n-1}\backslash B^{n-1}\to\mathbb{C}$ with
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\gtrsim R^{\varepsilon}{\left\|f\right\|}_{2}$ (3.26)
Here, $A\in\mathbb{N}$ is chosen to be sufficiently large.
Output. The algorithm produces the following objects:
  • A collection $\mathcal{O}$ of cells, each contained in a $R^{O(\delta_{0})}$-ball.
  • For each $O\in\mathcal{O}$, a function $f_{O}$.
  • Scales $\overrightarrow{r}=(r_{n},\dots,r_{m-1})$ satisfying $R=r_{n}>r_{n-1}>\cdots>r_{m-1}=1$.
  • Integers $\overrightarrow{A}=(A_{n},\dots,A_{m-1})$ satisfying $A=A_{n}>A_{n-1}>\cdots>A_{m-1}$.
  • A tuple of parameters $\overrightarrow{D}=(D_{n},\dots,D_{m-1})$.
  • For $m\leq l\leq n$, a family $\overrightarrow{\mathcal{S}}_{l}$ of level (nl+1)(n-l+1) multigrains. Each $\overrightarrow{S}_{l}=((Z_{l},B_{r_{l}}),\dots,(Z_{n},B_{r_{n}}))\in\overrightarrow{\mathcal{S}}_{l}$ has multiscale $\overrightarrow{r}_{l}=(r_{l},\dots,r_{n})$ and bounded degree. Moreover, for $l\leq s\leq t\leq n$,
    $Z_{s}=Z^{\prime}_{s}+b\quad\text{for some }b\text{ with }|b|\lesssim r_{t}^{1/2+\delta_{t}},\qquad Z^{\prime}_{s}\subseteq Z_{t}.$
  • For $m\leq l\leq n$, an assignment of a function $f_{\overrightarrow{S}_{l}}$ to each $\overrightarrow{S}_{l}\in\overrightarrow{\mathcal{S}}_{l}$. Each $f_{\overrightarrow{S}_{l}}$ is tangent to $S_{l}=(Z_{l},B_{r_{l}})$, the first component of $\overrightarrow{S}_{l}$.
Property:
1.The inequality
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox M\left(\overrightarrow{r},\overrightarrow{D}\right){\left\|f\right\|}^{1-{\beta}_{m}}_{2}{\left(\sum_{O\in\mathcal{O}}{{\left\|Ef_{O}\right\|}^{p_{m}}_{BL^{p_{m}}_{k,A_{m-1}}\left(O\right)}}\right)}^{{\beta}_{m}/p_{m}\ }$ (3.27)
holds, where
$M\left(\overrightarrow{r},\overrightarrow{D}\right):={D^{\delta}_{m-1}\left(\prod^{n-1}_{i=m}{D_{i}}\right)}^{\left(n-m\right)\delta}\left(\prod^{n-1}_{i=m}{r^{\left({\beta}_{i+1}-{\beta}_{i}\right)/2}_{i}D^{\left({\beta}_{i+1}-{\beta}_{m}\right)/2}_{i}}\right)$ (3.28)
2.The inequality
$\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\lessapprox\left(\prod^{n-1}_{i=m-1}{D_{i}}\right){\left\|f\right\|}^{2}_{2}$ (3.29)
holds.
3.For $m\leq l\leq n$,
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox r^{-\frac{n-l}{2}}_{l}\left(\prod^{l-1}_{i=m-1}{r^{-1/2}_{i}D^{-i}_{i}}\right){\mathop{\mathrm{max}}_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\ }$ (3.30)
[…]
${\left\|f_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox r^{\frac{n-l-1}{2}}_{l}\prod^{n-1}_{j=l+1}{r^{-1/2}_{j}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}$ (3.31)
where $f^{\#}_{{\overrightarrow{S}}_{l}}:={\left(f_{{\overrightarrow{S}}_{n-1}}\right)}_{\mathbb{T}\left[{\overrightarrow{S}}^{\#}_{l}\right]}$.
Initialization. We set
$r_{n}:=R,D_{n}:=1,A_{n}:=A$
${\overrightarrow{\mathcal{S}}}_{n}:=\left\{{\overrightarrow{S}}_{n}\right\},{\overrightarrow{S}}_{n}:=\left(S_{n}\right),S_{n}=\left({\mathbb{R}}^{n},\ B_{R}\right),f_{{\overrightarrow{S}}_{n}}:=f$
$\left(n+2-l\right)$-th step. Suppose the algorithm has run for (n+1l)(n+1-l) steps and the following things have been determined.
${\overrightarrow{r}}_{l}=\left(r_{n},\dots,r_{l}\right),\ \ {\overrightarrow{A}}_{l}=\left(A_{n},\dots,A_{l}\right),\ \ {\overrightarrow{D}}_{l}=\left(D_{n},\dots,D_{l}\right),\ \ {\overrightarrow{\mathcal{S}}}_{l}$
For each $\overrightarrow{S}_{l}\in\overrightarrow{\mathcal{S}}_{l}$, we apply Algorithm 1 to the $k$-broad norm $\|Ef_{\overrightarrow{S}_{l}}\|_{BL^{p_{l}}_{k,A_{l}}(B_{r_{l}})}$. Let $r[\overrightarrow{S}_{l}]$, $D[\overrightarrow{S}_{l}]$, and $A[\overrightarrow{S}_{l}]$ denote the quantities $r_{J}$, $D$, and $A_{J}$ returned by Algorithm 1, respectively. Let
${\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}:=\left\{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}\mathrm{:}\mathrm{[alg\ 1]\ terminates\ due\ to\ [t}\mathrm{iny}\mathrm{]}\right\}$
${\overrightarrow{\mathcal{S}}}_{l\mathrm{,tang}}:=\left\{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}\mathrm{:}\mathrm{[alg\ 1]\ terminates\ due\ to\ [t}\mathrm{ang}\mathrm{]}\right\}$
Consequently, at least one of the following two cases must hold:
Tiny-dominated Case. We say that the tiny-dominated case happens if
$\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\leq 2\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$ (3.32)
By (3.15), there exists ${\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}\subseteq{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}$ and $A_{l-1}\in\mathbb{N}$ such that $A\left[{\overrightarrow{S}}_{l}\right]=A_{l-1}$ and $D\left[{\overrightarrow{S}}_{l}\right]=D_{l-1}$ for all ${\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}$ and
$\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\lessapprox\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$ (3.33)
For each $\overrightarrow{S}_{l}\in\overrightarrow{\mathcal{S}}^{\prime}_{l,\mathrm{tiny}}$, let $\mathcal{O}[\overrightarrow{S}_{l}]$ denote the final cells produced by Algorithm 1. For each $O\in\mathcal{O}[\overrightarrow{S}_{l}]$, let $f_{O}$ be the corresponding function returned by Algorithm 1. Algorithm 2 then outputs the following additional data and terminates:
$m=l,\qquad r_{m-1}:=1,\qquad\mathcal{O}:=\bigcup_{\overrightarrow{S}_{m}\in\overrightarrow{\mathcal{S}}^{\prime}_{m,\mathrm{tiny}}}\mathcal{O}[\overrightarrow{S}_{m}].$
Tangent-dominated Case. If (3.32) fails, then
$\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\leq 2\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tang}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$
By (3.15), there exists ${\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tiny}}\subseteq{\overrightarrow{\mathcal{S}}}_{l\mathrm{,tiny}}$, $A_{l-1}\in\mathbb{N}$, and $r_{l-1}\in\left(0,R^{O\left({\delta}_{0}\right)}\right]$ such that
$\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}_{l}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}\lessapprox\sum_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}}{{\left\|Ef_{{\overrightarrow{S}}_{l}}\right\|}^{p_{l}}_{BL^{p_{l}}_{k,A_{l}}\left(B_{r_{l}}\right)}}$ (3.34)
and
$A\left[{\overrightarrow{S}}_{l}\right]=A_{l-1},\ \ r\left[{\overrightarrow{S}}_{l}\right]=r_{l-1},\ \ D\left[{\overrightarrow{S}}_{l}\right]=D_{l-1},\ \ \forall{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}$
For each $\overrightarrow{S}_{l}\in\overrightarrow{\mathcal{S}}^{\prime}_{l,\mathrm{tang}}$, let $S[\overrightarrow{S}_{l}]$ denote the collection of grains produced by Algorithm 1. For each $S_{l-1}\in S[\overrightarrow{S}_{l}]$, define
$\overrightarrow{S}_{l-1}:=(S_{l-1},\overrightarrow{S}_{l}),$
and let $f_{\overrightarrow{S}_{l-1}}$ denote the corresponding function returned by Algorithm 1. Finally, we define
${\overrightarrow{\mathcal{S}}}_{l\mathrm{-}\mathrm{1}}:=\bigcup_{{\overrightarrow{S}}_{l}\in{\overrightarrow{\mathcal{S}}}^{\prime}_{l\mathrm{,tang}}}{\left\{{\overrightarrow{S}}_{l-1}:S_{l\mathrm{-}\mathrm{1}}\in S\left[{\overrightarrow{S}}_{l}\right]\right\}}$
Verification of the Properties. Property 2 follows from (3.22) together with repeated applications of (3.17). Property 3 follows from (3.23) and repeated applications of (3.18). Property 1 follows from (3.16), (3.34), (3.21), (3.33), Lemmas 2.12 and 2.13; see [11] for details. We now verify Property 4.
Applying (3.20) repeatedly yields
$\|f_{\overrightarrow{S}_{l}}\|_{2}^{2}\lessapprox\prod_{i=l}^{n-2}\left(\frac{r_{i+1}}{r_{i}}\right)^{-(n-i-1)/2}\|f^{\#}_{\overrightarrow{S}_{l}}\|_{2}^{2}=r_{l}^{(n-l-1)/2}\prod_{j=l+1}^{n-1}r_{j}^{-1/2}\|f^{\#}_{\overrightarrow{S}_{l}}\|_{2}^{2}.$
This establishes Property 4.
3.3. Establish the Broad Estimates
We now apply the properties provided by Algorithm 2. Since each $O\in\mathcal{O}$ has diameter at most $R^{O(\delta_{0})}$, we may trivially bound
${\left\|Ef_{O}\right\|}_{BL^{p_{m}}_{k,A_{m-1}}\left(O\right)}\lessapprox{\left\|f_{O}\right\|}_{2}$
Together with (3.27), this yields
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox M\left(\overrightarrow{r},\overrightarrow{D}\right){\left\|f\right\|}^{1-{\beta}_{m}}_{2}{\left(\sum_{O\in\mathcal{O}}{{\left\|f_{O}\right\|}^{2}_{2}}\right)}^{{\beta}_{m}/p_{m}\ }{\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2{\beta}_{m}\left(\frac{1}{2}-\frac{1}{p_{m}}\right)}_{2}\ }$ (3.35)
Combining this with (3.28) and (3.25), we obtain
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{i=m-1}{r^{\frac{{\beta}_{i+1}-{\beta}_{i}}{2}}_{i}D^{\frac{{\beta}_{i+1}}{2}-\left(\frac{1}{2}-\frac{1}{p}\right)}_{i}}{\left\|f\right\|}^{\frac{2}{p}}_{2}{\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{1-\frac{2}{p}}_{2}\ }$ (3.36)
4. Complete the proof
In this section, we complete the proof of Proposition 2.16. For the reader’s convenience, we restate the proposition below.
Proposition 4.1.
Let $n\geq 3$ and $k\in[2,n-2]$. For
$p=2+\frac{12}{4n-5+2\left(k-1\right)\prod^{n-2}_{i=k}{\frac{2i}{2i+1}}},$
$\frac{1}{q}=\frac{1}{p}+\frac{1}{4n-8}\big(\frac{6}{p}-1\big),$
and some integer $A\lesssim_{\varepsilon,n}1$, the estimate
$\|Ef\|_{BL_{k,A}^{p}(B_{R})}\lesssim_{p,q,\varepsilon}R^{\varepsilon}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{\frac{2}{q}}\|f\|_{L^{\infty}(\mathbb{R}^{n-1})}^{1-\frac{2}{q}}$
holds uniformly for all Schwartz functions $f$ with $\operatorname{supp}f\subset 2\overline{B}^{\,n-1}\setminus B^{n-1}$ and $R\geq 1$.
Suppose that, after applying Algorithm 2 to decompose the $k$-broad norm $\|Ef\|_{BL^{p}_{k,A}(B_{R})}$, we obtain the outputs described in Section 3.2. By Lemma 2.14, we have $m\geq k$. Consequently, we divide the proof into cases $n-1\leq m\leq n$ and $k\leq m\leq n-2$.
Case $n-1\leq m\leq n$ . Taking $l=n$ in (3.30), we obtain
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\left(\prod^{n-1}_{j=m-1}{r^{-1/2}_{j}D^{-j}_{j}}\right){\left\|f\right\|}^{2}_{2}$ (4.1)
Combining this with (3.36), we have
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox\prod^{n-1}_{j=m-1}{r^{\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}D^{\frac{{\beta}_{j+1}}{2}-\left(j+1\right)\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}}{\left\|f\right\|}^{2}_{2}$ (4.2)
Set
$\begin{cases}p=p_{n}=\frac{2n}{n-1}&\text{if $m=n$},\\[4.0pt] p_{n-1}=\frac{2(n-1)}{n-2},\quad p=p_{n}=\frac{2(2n-1)}{2n-3}&\text{if $m=n-1$}.\end{cases}$ (4.3)
In both cases, $p<p(n,k)$ and
$\frac{{\beta}_{i+1}-{\beta}_{i}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)=0,\quad\forall j\in[m-1,n-1]$
$\frac{{\beta}_{i+1}}{2}-\left(i+1\right)\left(\frac{1}{2}-\frac{1}{p}\right)\leq 0,\quad\forall j\in[m-1,n-1]$
This finishes the proof in this case.
Case $k\leq m\leq n-2$ . Applying the nested polynomial Wolff axioms, we have
Lemma 4.2.
For $m\leq l\leq n-2$,
${\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lessapprox\left(\prod^{n-2}_{i=l}{r^{-\frac{1}{2}}_{i}}\right){\left\|f\right\|}^{2}_{\infty}$
Proof.
Since $f^{\#}_{\overrightarrow{S}_{l}}$ is concentrated on the $r_{n-1}$-scale wave packets $(\theta,v,l)$ with $\theta\in\Theta[\overrightarrow{S}^{\#}_{l}$], Corollary 2.27 implies
${\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{2}\lesssim\#\mathrm{\Theta}\left[{\overrightarrow{S}}^{\#}_{l}\right]r^{-\frac{n-2}{2}}_{n-1}{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\\ \lessapprox\left(\prod^{n-2}_{j=l}{r^{-\frac{1}{2}}_{j}}\right){\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (4.4)
By Corollary 2.3,
${\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lesssim{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f_{{\overrightarrow{S}}_{n-1}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$
Together with (3.19) we have
${\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f^{\#}_{{\overrightarrow{S}}_{l}}\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }\lessapprox{\mathop{\mathrm{max}}_{\theta:r^{-1/2}_{n-1}\mathrm{-}\mathrm{sector}}{\left\|f\right\|}^{2}_{L^{2}_{\mathrm{avg}}\left(\theta\right)}\ }$ (4.5)
Lemma (4.2) now follows from (4.4) and (4.5). ∎
Combining (3.30), (3.31), and Lemma 4.2, we obtain that for $m\leq l\leq n-2$,
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\left(\prod^{n-1}_{j=m\ }{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{n-2}_{j=l}{r^{-\frac{1}{2}}_{j}}\right)\left(\prod^{l-1}_{j=m-1}{D^{-j}_{j}}\right){\left\|f\right\|}^{2}_{\infty}$ (4.6)
Define $p_{j}$ by
$p_{j}:=\begin{cases}2+\frac{6}{2\left(j-1\right)+\left(m-1\right)\prod^{j-1}_{i=m}{\frac{2i}{2i+1}}}&\text{if $m\leq j\leq n-1$},\\[4.0pt] p(n,k)&\text{if $j=n$}.\end{cases}$ (4.7)
For $m\leq j\leq n-1$, we define $\gamma_{j}\in[0,1]$ by
$\gamma_{j}:=\begin{cases}\frac{m-1}{2}\frac{1}{j\left(j-1\right)}\prod^{j}_{i=m}{\frac{2i}{2i+1}}&\text{if $m\leq j\leq n-2$},\\[4.0pt] 1-\sigma_{n-2}&\text{if $j=n-1$}.\end{cases}$ (4.8)
where
${\sigma}_{j}:=\sum^{j}_{i=m}{{\gamma}_{i}}$
Taking the geometric mean of (4.6) with weights $\gamma_{l}$ and (4.1) with weight $\gamma_{n-1}$, we obtain
${\mathop{\mathrm{max}}_{O\in\mathcal{O}}{\left\|f_{O}\right\|}^{2}_{2}\ }\lessapprox\\ {r_{n-1}^{-\frac{1}{2}}}D_{n-1}^{-\gamma_{n-1}(n-1)}\left(\prod^{n-2}_{j=m}{r^{-\frac{1+{\sigma}_{j}}{2}}_{j}}\right)\left(\prod^{n-2}_{j=m-1}{D^{-j\left(1-{\sigma}_{j}\right)}_{j}}\right){\left\|f\right\|}^{2\gamma_{n-1}}_{2}{\left\|f\right\|}^{2\sigma_{n-2}}_{\infty}$
Together with (3.36), we see that
${\left\|Ef\right\|}_{BL^{p}_{k,A}\left(B_{R}\right)}\lessapprox{r^{\frac{{\beta}_{n}-{\beta}_{n-1}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{n-1}}D^{\frac{{\beta}_{n}}{2}-\left(1+\gamma_{n-1}(n-1)\right)\left(\frac{1}{2}-\frac{1}{p}\right)}_{n-1}\\ \prod^{n-2}_{j=m}{r^{\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1+{\sigma}_{j}}{2}\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}}\prod^{n-2}_{j=m-1}{D^{\frac{{\beta}_{j+1}}{2}-\left(1+j\left(1-{\sigma}_{j}\right)\right)\left(\frac{1}{2}-\frac{1}{p}\right)}_{j}}{\left\|f\right\|}^{\frac{2}{q}}_{2}{\left\|f\right\|}^{1-\frac{2}{q}}_{\infty}$
where $q$ is defined by
$\frac{1}{q}=\frac{1}{p}+\frac{1}{4n-8}\big(\frac{6}{p}-1\big)$
Arguing as in [12], we see that
$\frac{{\beta}_{n}-{\beta}_{n-1}}{2}-\frac{1}{2}\left(\frac{1}{2}-\frac{1}{p}\right)=0$
$\frac{{\beta}_{n}}{2}-\left(1+\gamma_{n-1}(n-1)\right)\left(\frac{1}{2}-\frac{1}{p}\right)<0$
$\frac{{\beta}_{j+1}-{\beta}_{j}}{2}-\frac{1+{\sigma}_{j}}{2}\left(\frac{1}{2}-\frac{1}{p}\right)=0,\qquad\forall j\in[m,n-2]$
$\frac{{\beta}_{j+1}}{2}-\left(1+j\left(1-{\sigma}_{j}\right)\right)\left(\frac{1}{2}-\frac{1}{p}\right)=0,\qquad\forall j\in[m-1,n-2]$
This concludes the proof since $k\leq m$.